### v3

# A Volumetric Approach to Privacy of Dynamical Systems

Abstract
Information-theoretic metrics, such as mutual information, have been widely used to evaluate privacy leakage in dynamic systems. However, these approaches are typically limited to stochastic systems and face computational challenges. In this paper, we introduce a novel volumetric framework for analyzing privacy in systems affected by unknown but bounded noise. Our model considers a dynamic system comprising public and private states, where an observation set of the public state is released. An adversary utilizes the observed public state to infer an uncertainty set of the private state, referred to as the inference attack. We define the evolution dynamics of these inference attacks and quantify the privacy level of the private state using the volume of its uncertainty sets. We then develop an approximate computation method leveraging interval analysis to compute the private state set. We investigate the properties of the proposed volumetric privacy measure and demonstrate that it is bounded by the information gain derived from the observation set. Furthermore, we propose an optimization approach to designing privacy filter using randomization and linear programming based on the proposed privacy measure. The effectiveness of the optimal privacy filter design is evaluated through a production-inventory case study, illustrating its robustness against inference attacks and its superiority compared to a truncated Gaussian mechanism.
keywords:
Volumetric privacy measure; privacy protection; interval analysis; truncated Gaussian mechanism.
thanks: The work was partially supported by the Research Grants Council of Hong Kong under Project CityU 21208921, a grant from Chow Sang Sang Group Research Fund sponsored by Chow Sang Sang Holdings International Limited.
1 Introduction
1.1 Motivation
Data sharing plays a pivotal role in enabling cooperative decision-making and optimization in dynamic processes. However, the exposure of such data may inadvertently reveal sensitive information. Specifically, correlations between shared metrics and underlying operational parameters can be exploited by adversaries to develop competitive and malicious strageties. This challenge highlights the critical need for methodologies that preserve data utility while ensuring rigorous privacy protection for dynamic systems.
Information-theoretic metrics, including mutual information, directed information, and conditional entropy, have been extended to quantify privacy leakage in dynamic systems. Despite their utility, these metrics have notable limitations. For instance, as highlighted in [8, 30], mutual information evaluates privacy leakage in an average sense, which can overlook privacy breaches for infrequent realizations of private variables. The computational complexity of these measures is another challenge, as they require evaluating leakage for all possible realizations of random variables [25, 32, 11]. Most existing information-theoretic approaches are focused on stochastic systems and typically assume complete knowledge of the probability distributions of system states. However, this assumption does not hold in systems that are influenced by unknown but bounded (UBB) noise.
To address these challenges, this paper introduces a volumetric privacy measure based on the uncertainty set of the private variable. The proposed approach is applicable to both deterministic and stochastic systems and eliminates the need for prior knowledge of probability distributions.
1.2 Contributions
This paper addresses the privacy protection problem for dynamic systems with UBB noise, as illustrated in Fig. 1. In this framework, the system state is partitioned into two categories: the public state $X_{k}$ and the private state $Y_{k}$, both of which belong to a known bounded set. The observation of the system states is represented by the set $\mathcal{M}_{k|k}^{x}$ which contains the actual public state $X_{k}=x_{k}$. An untrusted third party, i.e., the adversary, uses $\mathcal{M}_{k|k}^{x}$ to infer the private state by constructing an uncertainty set $\mathcal{Y}_{k|k}$, which is referred to as the inference attack.
Figure 1: The inference attack.
This system setup is inspired by the production-inventory problem discussed in Section 2. In such scenarios, inventory information can aid distributors in optimizing sales. However, it also enables adversaries to infer sensitive information, such as the production rate, which reveals a company’s private data, including production efficiency, strategies, and supply chain dynamics.
The contributions of this paper are summarized as follows:
  • Definition of Inference Attack: We define the adversary’s inference attack using set operations.
  • Volumetric Privacy Measure: A novel privacy measure is proposed based on the volume of the uncertainty set for the private state.
  • Privacy Level Computation: We develop a computational method utilizing interval analysis to evaluate the privacy level.
  • Optimal Privacy Filter Design: An optimization-based approach is introduced to design privacy filters, based on randomization and linear programming.
Additionally, we validate our approach using the production-inventory example. The results demonstrate that the adversary can infer sensitive production rates, through inventory data. However, this leakage can be significantly mitigated under the proposed optimal privacy filter design. Also, compared with the truncated Gaussian mechanism, our privacy filter design achieves lower data distortion while maintaining the same privacy level.
1.3 Related work
Differential privacy was initially defined to hide the presence of an individual or a specific record in the responses of queries for static datasets. The authors in [15] extended the differential privacy to dynamic systems, and proposed the differentially private Kalman filter for linear systems. In [5], the authors proposed a general architecture for private filtering by approximating the desire filter in a differential private way, and studied the application of differential privacy in distributed control and optimization. In [21], the authors studied the differential privacy preserving average consensus problem to protect the initial state in distributed control by adding and abstracting random noises. The authors in [12] connected the input observability with differential privacy and proposed the privacy preserving controller design method. The authors in [33] studied the minimal amount noises added to multi-agent systems for differential privacy based on the minimal observability subspace. They proposed node-based and edge-based privacy-preserving mechanisms based on the optimization of the minimum added noise, and discussed the trade-offs between utility and privacy. In [34], the authors proposed several cost-friendly differential privacy-preserving schemes by randomly adjusting the charge-discharge rate of a batter based on a modified Laplace distribution.
Information-theoretic protection methods use conditional entropy, mutual information, etc, from information theory, to measure privacy leakage. In [22], the authors studied the privacy filter design for a Markov chain to hide the private state when transmitting the monitored state to a receiver, which was approximately solved with a greedy algorithm based on the convex optimization. The authors in [3] designed a private filter to protect the state of a hidden Markov chain via minimizing the mutual information with the constraint of the observation utility. In [31, 24], the authors studied the privacy-aware filter design for linear systems, using the directed information between the system state and the output of the filter as the privacy measure. It was proved that the optimal filter can be realized with a Kalman filter and an additive Gaussian mechanism.
The authors in [32] studied the structural properties of the privacy-aware state estimation problem based on the dynamic programming decomposition. They approximated the privacy leakage with variational techniques and solved the estimation problem via policy gradient approaches. In [23], the conditional entropy of the private state given measurements and controls was used as a penalty term in the partially observable Markov decision process (POMDP) optimization, to avoid the adversary accurately estimate the system state. They showed that the state obfuscation problem is a standard POMDP optimization problem in which the cost-to-go function is concave in the belief state. In [16], the authors considered to protect the privacy of the electricity demand of a household to avoid the adversary infers the indoor activity. They utilized the rechargeable battery to hide the actual electricity demand and obtained the optimal charging policy via minimizing the privacy leakage measured by mutual information. However, the classical information-theoretic metrics for privacy might have some limitations when facing one-try attack as discussed in [30], which motivates the development of $R\acute{e}nyi$ min-entropy [30], maximal leakage [9] and pointwise maximal leakage [28], etc.
In most existing privacy protection literature for dynamic systems, it is assumed that the system are disturbed with unbounded noise that has probabilistic properties, e.g., Gaussian distribution. However, there are also dynamic systems are driven with UBB noise without probability assumptions. The privacy concept for these systems are not well developed. The authors in [6, 7] designed differentially private set-based estimator to protect the privacy of measurements based on a truncated noise distribution. In [13], the authors proposed the guaranteed privacy extent from differential privacy to hide the measurement sensor identities, and provided the optimization method for the $\mathcal{H}_{\infty}$ optimal privacy preserving interval observer design. Furthermore, in [26, 27], authors studied the state opacity problems where the discrete state space is divided into disjoint sets of secrete and non-secrete states, and the system outputs from secrete and non-secrete sets are indistinguishable. The authors in [19] proposed $\delta-$approximate initial-state opacity as a secrete notion for discrete-time systems with continuous states, and developed verification approaches based on Barrier certificates.
The proposed volumetric privacy measure in this work is inspired by set-membership estimation methods for systems with UBB noise [20]. Most set-membership methods represent states as bounded geometric sets, such as intervals [10], zonotopes [14], or ellipses [4], where the volumes of these sets correspond to estimation uncertainty. Building on these methods, we consider the adversary’s ability to infer the private state set based on observations of the public state. Consequently, a larger volume of the uncertainty set implies a broader range of potential private states that could be consistent with the observed public state set. This motivates our use of the volume of the private state set as a privacy measure.
This work differs from prior studies, such as [6], [7], and [13], which focus on ensuring the differential privacy of system states. In contrast, we address the problem of protecting private states while publicly releasing non-sensitive states, as illustrated in Fig. 1. Our approach uses a volumetric privacy measure, which also distinguishes it from state opacity problems [26], [27], and [19], where the state space is partitioned into secret and non-secret sets.
Furthermore, state opacity problems focus on guaranteeing privacy by ensuring that at least one non-secret state produces an observation indistinguishable from the secret state. In these frameworks, enlarging the volume of the non-secret state set does not enhance privacy, as privacy is determined solely by the existence of indistinguishable outputs, not by the size of the state set. Differently, within our framework, a larger private state set is preferred, as it increases the adversary’s uncertainty in inferring the private state.
1.4 Outline
The rest of the paper is organized as follows. Section 2 introduces the system model and the inference attack. Section 3 defines the volumetric privacy measure, provides computational approaches for privacy level evaluation, and discusses the properties of the proposed measure. Section 4 presents an optimal privacy filter design to mitigate privacy leakage while maintaining a certain utility level. Section 5 presents numerical results, followed by the conclusions in Section 6.
1.5 Notation
We use italic letters to denote the set of unknown variables, e.g., $\mathcal{X}$ and $\mathcal{Y}$ for $X$ and $Y$. We use the vector $\left[\begin{array}[]{c}\underline{X}\\ \overline{X}\\ \end{array}\right]$ to describe the interval $\mathcal{X}$, i.e., $\mathcal{X}=\left\{X|\underline{X}\leqslant X\leqslant\overline{X}\right\}$, where $\underline{X}$ and $\overline{X}$ are the lower and upper endpoints, respectively. The interval $\mathcal{X}$ can also be represented with
$\mathcal{X}=\left\{c^{x}+\mathrm{diag}\left(p^{x}\right)\alpha:\alpha\in\mathcal{R}^{n_{x}},\left|\alpha\right|_{\infty}\leqslant 1\right\},$
where $c^{x}=\frac{\overline{X}+\underline{X}}{2}$ and $p^{x}=\frac{\overline{X}-\underline{X}}{2}$ are the center point and radius of the interval. Besides, given the block matrix $A=\left[A_{1},A_{2}\right]$, the multiplication between the matrix $A$ and the interval $\mathcal{X}$ is defined as $A\mathcal{X}=A_{1}\underline{X}+A_{2}\overline{X}$. Also, if $\mathcal{X}$ and $\mathcal{Y}$ are intervals, then we use $\mathcal{X}+\mathcal{Y}$ to denote the interval $\left[\begin{array}[]{c}\underline{X}+\underline{Y}\\ \overline{X}+\overline{Y}\\ \end{array}\right]$, and use $\mathcal{X}-\mathcal{Y}$ to represent the interval $\left[\begin{array}[]{c}\underline{X}-\overline{Y}\\ \overline{X}-\underline{Y}\\ \end{array}\right]$. As for the non-interval set $\mathcal{Z}$, we use $A\mathcal{Z}$ to denote the new set $\left\{AZ|Z\in\mathcal{Z}\right\}$, and use $\mathcal{Z}\oplus\mathcal{R}$ to represent $\left\{Z+R|X\in\mathcal{Z},R\in\mathcal{R}\right\}$. Furthermore, the 1-norm of the column vector $b$ with n dimensions is defined as $\left\|b\right\|_{1}=\sum_{i=1}^{n}\left|b\left(i\right)\right|$, and $b^{\top}$ is the transpose of $b$. The vector $\mathbf{1}_{n_{x}}$ denotes a column vector of ones with $n_{x}$ dimensions, while $I_{n_{x}\times n_{x}}$ represents an identity matrix of size $n_{x}\times n_{x}$.
2 System Model and Inference Attack
2.1 System Model
We consider the following stable system model $\mathbf{G_{1}}$,
$\mathbf{G_{1}}:\left\{\begin{array}[]{c}X_{k}=A_{1}X_{k-1}+A_{2}Y_{k-1}+B_{1}W_{k}^{x}\\ Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+B_{2}W_{k}^{y}\\ \end{array}\right.,$ (3)
where $A_{1}$ and $A_{2}$ are invertible, $Y_{k}\in\mathcal{R}^{n}$ is the private state, $X_{k}\in\mathcal{R}^{n}$ is the public state to be released, $W_{k}^{x}\in\mathcal{W}_{k}^{x}\subseteq\mathcal{R}^{m}$ and $W_{k}^{y}\in\mathcal{W}_{k}^{y}\subseteq\mathcal{R}^{m}$ are the UBB noises. Also, the initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively. When $A_{3}$ is zero, the model $\mathbf{G_{1}}$ can be regarded as a dynamic system with the public state $X_{k}$ and the private input $Y_{k}$.
For analysis convenience, we assume that $\mathcal{X}_{0|-1}$, $\mathcal{Y}_{0|-1}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$ are intervals. For other types of bounded sets, such as zonotopes [14] and ellipses [4], the analysis is left for future investigation. Note that we do not assume the prior knowledge of probability distributions of the process noises, e.g., Gaussian distribution, which is different from most existing literature in privacy protection for dynamic systems. Furthermore, we assume that the adversary has full knowledge of system model $\mathbf{G_{1}}$ and will collect information of the public state to infer the private state.
2.2 Motivating Example
Consider the problem of privacy leakage related to the production rate in inventory control systems. To enhance sales, companies often need to share inventory levels with distributors. However, inventory levels are closely correlated with the production rate, which serves as a key indicator of the factory’s production strategy, operational efficiency, and supply chain dynamics. Therefore, the production rate constitutes highly sensitive information. If accessed by an adversary, this data could be exploited to develop competitive strategies, potentially compromising the company’s market position and operational integrity.
The inventory of production can be modeled as follows [17, 29],
$X_{k}=X_{k-1}+Y_{k-1}-W_{k}^{x},$ (4)
where $X_{k}$ is the inventory, $Y_{k}$ is the production rate, and $W_{k}^{x}$ is the uncertain demand. The production rate is usually chosen to maintain the inventory around a desired level, which can be modeled as follows,
$Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+W_{k}^{y},$ (5)
where the uncertain process noise $W_{k}^{y}$ is due to the adversary’s limited knowledge about the company’s production capability.
When the uncertainty of demand and production rate is small, e.g., $W_{k}^{x}=D^{x}$ and $W_{k}^{y}=D^{y}$ are constant, then the adversary can approximately reconstruct the production rate with the inventory level via
$Y_{k-1}\approx X_{k}-X_{k-1}+D^{x},$ (6)
and
$Y_{k}\approx A_{3}X_{k-1}+A_{4}Y_{k-1}+D^{y},$ (7)
which causes privacy leakage and increases potential risks of business competition.
However, in practical scenarios, the uncertainty associated with demand and production capability is bounded but cannot be ignored. Inferring the private production rate, therefore, becomes a complex but still feasible task. Due to the presence of unknown noise terms that belong to bounded sets without probabilistic assumptions, multiple private states may correspond to the same public state with equal probability, leading to the formation of an uncertainty set. To address this challenge, we define the adversary’s inference attack on the private state using set operations in the following sections.
2.3 Inference Attack
We assume that the adversary observes a set of public states, denoted as $\mathcal{M}_{k|k}^{x}$, which includes the actual public state value $X_{k}=x_{k}$. Also, every element of $\mathcal{M}_{k|k}^{x}$ is considered a potential candidate for the actual public state. Notably, if the public state is directly transmitted to the adversary without any processing, the public state set $\mathcal{M}_{k|k}^{x}$ contains only the actual public state.
Based on the observed public state set, the adversary identifies all potential values of the private state that align with $\mathcal{M}_{k|k}^{x}$ to construct its uncertainty set. This process is referred to as the inference attack. We next define the inference attack recursively.
At time $k$, given the public state set $\mathcal{X}_{k-1|k-1}$ and the uncertainty private state set $\mathcal{Y}_{k-1|k-1}$, the set of public state can be predicted based on the system model (3), i.e.,
$\mathcal{X}_{k|k-1}=A_{1}\mathcal{X}_{k-1|k-1}\oplus A_{2}\mathcal{Y}_{k-1|k-1}\oplus B_{1}\mathcal{W}_{k}^{x}.$ (8)
Similarly, the set of private state can be predicted via
$\mathcal{Y}_{k|k-1}=A_{3}\mathcal{X}_{k-1|k-1}\oplus A_{4}\mathcal{Y}_{k-1|k-1}\oplus B_{2}\mathcal{W}_{k}^{y}.$ (9)
After receiving the observation set of the public state $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$, the adversary extracts new information from $\mathcal{M}^{x}_{k|k}$ and updates the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via the following steps,
$\mathcal{M}_{k-1|k}^{x}=$$\left\{M^{x}_{k-1|k}|X_{k|k}=A_{1}M^{x}_{k-1|k}+A_{2}Y_{k-1|k-1}+B_{1}W_{k}^{x},\right.$
$\forall X_{k|k}$ $\in\mathcal{M}^{x}_{k|k},\left.\forall Y_{k-1|k-1}\in\mathcal{Y}_{k-1|k-1},\forall W_{k}^{x}\in\mathcal{W}_{k}^{x}\right\}\!,$ (10)
$\mathcal{M}_{k-1|k}^{y}=$$\left\{M^{y}_{k-1|k}|X_{k|k}=A_{1}X_{k-1|k}+A_{2}M^{y}_{k-1|k}+B_{1}W_{k}^{x},\right.$
$\forall X_{k|k}$ $\in\mathcal{M}^{x}_{k|k},\left.\forall X_{k-1|k-1}\in\mathcal{X}_{k-1|k-1},\forall W_{k}^{x}\in\mathcal{W}_{k}^{x}\right\}\!,$ (11)
$\mathcal{X}_{k-1|k}=\mathcal{M}_{k-1|k}^{x}\cap\mathcal{X}_{k-1|k-1},$ (12)
$\mathcal{Y}_{k-1|k}=\mathcal{M}_{k-1|k}^{y}\cap\mathcal{Y}_{k-1|k-1},$ (13)
where it first computes the possible sets of the public and private states, i.e., $\mathcal{M}_{k-1|k}^{x}$ and $\mathcal{M}_{k-1|k}^{y}$, based on the system model (3) and the observation $\mathcal{M}^{x}_{k|k}$ in (2.3) and (2.3), and then reduces the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via intersection operations in (12) and (13).
According to the system dynamics (3), the adversary estimates the uncertainty set of $Y_{k}$ via the following forward inference,
$\mathcal{Y}_{k|k}=A_{3}\mathcal{X}_{k-1|k}\oplus A_{4}\mathcal{Y}_{k-1|k}\oplus B_{2}\mathcal{W}_{k}^{y}.$ (14)
Finally, the public state set can be further calibrated via intersection,
$\mathcal{X}_{k|k}=\mathcal{M}^{x}_{k|k}\cap\mathcal{M}^{x}_{k|k-1},$ (15)
where
$\mathcal{M}^{x}_{k|k-1}=A_{1}\mathcal{X}_{k-1|k}\oplus A_{2}\mathcal{Y}_{k-1|k}\oplus B_{1}\mathcal{W}_{k}^{x},$ (16)
is the predicted uncertainty set of $X_{k}$ based on the calibrated sets $\mathcal{X}_{k-1|k}$ and $\mathcal{Y}_{k-1|k}$.
Starting from $k=0$, with the initial uncertainty sets $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, the adversary can recursively update the uncertainty sets of $X_{k}$ and $Y_{k}$ via the backward calibration (2.3)-(13), and the forward inference (14)-(16). The backward calibration (2.3)-(13) reduces the uncertainty of $X_{k-1}$ and $Y_{k-1}$, which leads to the following proposition.
Proposition 1.
For any $k\geqslant 1$, the adversary’s uncertainty private state set (14) is a subset of its prediction set (9). Also, given the uncertainty sets $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$ containing the actual system states $X_{k-1}=x_{k-1}$ and $Y_{k-1}=y_{k-1}$, if the observation set $\mathcal{M}^{x}_{k|k}$ contains the actual public state $X_{k}=x_{k}$, then the inference result $\mathcal{Y}_{k|k}$ also contains the actual private state $Y_{k}=y_{k}$.
Proof 2.1.
The proof is omitted due to the simplicity.
According to Proposition 1, the inference attack results in a smaller uncertainty private state set $\mathcal{Y}_{k|k}$ that contains the actual private state $y_{k}$, meaning that there are fewer possible private states corresponds to the same observation set $\mathcal{M}_{k}^{x}$. In particular, if the uncertainty set $\mathcal{Y}_{k|k}$ contains only one element, then the adversary can obtain actual private state. In conclusion, the adversary can reduce its uncertainty of the private state via the inference attack.
2.4 Privacy Measure and Defense
The aforementioned inference attack relies on set operations, and its outcome is the uncertainty set of the private state, denoted as $\mathcal{Y}_{k|k}$. The first key question addressed in this paper is how to evaluate the privacy level of the private state using the uncertainty set $\mathcal{Y}_{k|k}$. Also, set operations, such as the addition in (2.3) and the intersection in (12), involve all possible combinations of elements from different collections. These operations are computationally challenging due to the continuous nature of the state space. Consequently, it is essential to provide efficient computational tools for privacy level evaluation. Moreover, from a defense perspective, another critical task is the design of effective mechanisms to mitigate privacy leakage.
To address these challenges, we propose a volumetric privacy measure and its computation approach in Section 3. Furthermore, we present an optimal privacy filter design to reduce privacy leakage in Section 4.
3 Volumetric Privacy Measure
In this section, we define the volume of uncertainty sets as a quantitative measure of the privacy level. We then introduce an interval-based approach to approximate the computation of state uncertainty sets and evaluate the corresponding privacy levels. Additionally, we analyze the properties of the volumetric measure, showing that both the privacy level and the uncertainty reduction are bounded by the new information derived from the observation set.
3.1 Privacy and Utility Measures
As discussed in Section 2.3, the uncertainty set of private state $\mathcal{Y}_{k|k}$ encompasses all possible elements that correspond to the same observation set, $\mathcal{M}_{k|k}^{x}$. However, due to the continuous nature of the state space, the number of elements in such sets is uncountable. To address this, we propose using the volume of the uncertainty set as a quantitative measure of privacy, defined as follows,
$P_{k}\left(\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right),$ (17)
where $\mathrm{Vol}\left(\cdot\right)\geqslant 0$ is a Lebesgue measure. When $\mathcal{Y}_{k|k}$ is an interval, we can compute its volume via $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)=\sum_{i=1}^{n}{\left(\overline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k}\left(i\right)\right)}$, where $\overline{Y}_{k|k}$ and $\underline{Y}_{k|k}$ are the upper and lower endpoints of the interval $\mathcal{Y}_{k|k}$, respectively.
Since the uncertainty set $\mathcal{Y}_{k|k}$ contains the actual value of private state $Y_{k}=y_{k}$, specially, the adversary can accurately access to $y_{k}$ if the volume of $\mathcal{Y}_{k|k}$ is zero. On the other hand, if the volume $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is large, then the range of private state will be large, which increases the difficulties of selecting the correct value as the estimation result. Therefore, the volume of $\mathcal{Y}_{k|k}$ describes the amount of the adversary’s inference uncertainty about the private state as well as the privacy level. We can increase the value of $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ to increase the privacy level to protect the system from inference attack.
Besides, since the uncertainty set $\mathcal{Y}_{k|k}$ is a subset of its predicted version $\mathcal{Y}_{k|k-1}$, we have
$\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\leqslant\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right),$ (18)
meaning that the adversary’s inference uncertainty about the private state is bounded by its prior knowledge $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$. In other words, the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ cannot exceed $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$.
Furthermore, we propose to use the following measure to measure the utility of public state set $\mathcal{X}_{k|k}$,
$U_{k}\left(\mathcal{X}_{k|k}\right)=\frac{1}{\mathrm{Vol}\left(\mathcal{X}_{k|k}\right)},$ (19)
since a larger $\mathcal{X}_{k|k}$ provides more possible values of the public state to the receiver, i.e., increases the estimation error. Consequently, a larger value of $U_{k}\left(\mathcal{X}_{k|k}\right)$ indicates a higher data utility of the public state set.
3.2 Inference Attack Approximation
In this subsection, we investigate the inference approximation for multi-dimensional systems using the tightest interval method [1]. This approach simplifies the computation of the privacy measure and provides the foundation for designing the optimal privacy filter as discussed in Section 4.
Lemma 2.
The tightest recursive inference interval from (2.3) to (13) can be computed via
$\mathcal{M}_{k-1|k}^{x}=$ $\Psi\left(A_{1}^{-1}\right)\mathcal{M}^{x}_{k|k}+\Psi\left(-A_{1}^{-1}A_{2}\right)\mathcal{Y}_{k-1|k-1}$
$+\Psi\left(-A_{1}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (20)
$\mathcal{M}_{k-1|k}^{y}=$ $\Psi\left(A_{2}^{-1}\right)\mathcal{M}^{x}_{k|k}+\Psi\left(-A_{2}^{-1}A_{1}\right)\mathcal{Y}_{k-1|k-1}$
$+\Psi\left(-A_{2}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (21)
$\mathcal{X}_{k-1|k}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k-1|k}^{x},\underline{X}_{k-1|k-1}\right\}\\ \min\left\{\overline{M}_{k-1|k}^{x},\overline{X}_{k-1|k-1}\right\}\\ \end{array}\right],$ (24)
$\mathcal{Y}_{k-1|k}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k-1|k}^{y},\underline{Y}_{k-1|k-1}\right\}\\ \min\left\{\overline{M}_{k-1|k}^{y},\overline{Y}_{k-1|k-1}\right\}\\ \end{array}\right],$ (27)
$\!\!\!\!\!\!\mathcal{M}_{k|k-1}^{x}\!=\!\Psi\!\left(A_{1}\right)\!\mathcal{X}_{k-1|k}\!+\!\Psi\!\left(A_{2}\right)\!\mathcal{Y}_{k-1|k}\!+\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x},$ (28)
$\mathcal{X}_{k|k}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k|k}^{x},\underline{M}_{k|k-1}^{x}\right\}\\ \min\left\{\overline{M}_{k|k}^{x},\overline{M}_{k|k-1}^{x}\right\}\\ \end{array}\right],$ (31)
$\!\!\!\!\!\!\!\!\mathcal{Y}_{k|k}=\Psi\left(A_{3}\right)\mathcal{X}_{k-1|k}+\Psi\left(A_{4}\right)\mathcal{Y}_{k-1|k}+\Psi\left(B_{2}\right)\mathcal{W}_{k}^{y},$ (32)
with
$\Psi\left(\star\right)=\left[\begin{matrix}\frac{\star+\left|\star\right|}{2}&\frac{\star-\left|\star\right|}{2}\\ \frac{\star-\left|\star\right|}{2}&\frac{\star+\left|\star\right|}{2}\\ \end{matrix}\right].$
Also, the tightest prior inference set of $Y_{k}$ is
$\!\!\!\!\!\!\!\mathcal{Y}_{k|k-1}\!=\!\Psi\left(\!A_{3}\!\right)\!\mathcal{X}_{k-1|k-1}\!+\!\Psi\left(\!A_{4}\!\right)\!\mathcal{Y}_{k-1|k-1}\!+\!\Psi\left(\!B_{2}\!\right)\!\mathcal{W}_{k}^{y},$ (33)
if $k\geqslant 1$. If $k=0$, then $\mathcal{Y}_{0|0}=\mathcal{Y}_{0|-1}$ and
$\mathcal{X}_{0|0}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{0|0}^{x},\underline{X}_{0|-1}\right\}\\ \min\left\{\overline{M}_{0|0}^{x},\overline{X}_{0|-1}\right\}\\ \end{array}\right].$ (36)
Proof 3.1.
See Appendix A.
According to Lemma 2, the inference attack can be approximately computed using intervals. It can be verified that the intervals computed in Lemma 2 are tight when $X_{k}$ and $Y_{k}$ are scalar variables.
3.3 Properties of the Inference Attack
The inference attack exhibits certain properties. For instance, the radius of the uncertainty $\mathcal{Y}_{k|k}$, i.e., $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$ is bounded by a function of the radius of the process noise and the observation set, as stated below.
Lemma 3.
For any $k\geqslant 1$, the radius of $\mathcal{Y}_{k|k}$ satisfies
$p_{k|k}^{y}\leqslant$ $\left(\left|A_{3}\right|+\left|A_{4}\right|\left|A_{2}^{-1}\right|+\left|A_{4}\right|\left|A_{2}^{-1}A_{1}\right|\right)\overline{p}^{x}$
$+\left|A_{4}\right|\left|A_{2}^{-1}B_{1}\right|p^{w,x}_{k}+\left|B_{2}\right|p^{w,y}_{k},$ (37)
where $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$ and $p^{w,y}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$, respectively.
Proof 3.2.
See Appendix B.
Since the volume of $\mathcal{Y}_{k|k}$ is the sum of $p_{k|k}^{y}$, the volume $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is also bounded by a function of $\overline{p}^{x}$. As a result, if the value of elements in $\overline{p}^{x}$ are small, i.e., if the observation set $\mathcal{M}_{k|k}^{x}$ is small, then the privacy level is low, and the adversary experiences less uncertainty after performing the inference attack.
Furthermore, by comparing the predicted and posterior uncertainty sets, as given by (9) and (14), the amount of uncertainty reduction can be quantified by $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, where
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\diagdown\mathcal{Y}_{k|k}.$
Proposition 4.
If $\mathcal{Y}_{k|k}$ is a subset of $\mathcal{Y}_{k|k-1}$, then the volume of $\Delta\mathcal{Y}_{k|k}$ is
$\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\!.$
Proof 3.3.
Since $\mathcal{Y}_{k|k}$ is the subset of $\mathcal{Y}_{k|k-1}$, the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is equal to $\left\|\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}\right\|_{1}$, which can be computed as follows,
$\left\|\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}\right\|_{1}$
$=$ $\sum_{i=1}^{n_{y}}\left(\overline{Y}_{k|k-1}\left(i\right)-\overline{Y}_{k|k}\left(i\right)+\underline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)$
$=$ $\sum_{i=1}^{n_{y}}{\left(\overline{Y}_{k|k-1}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)}-\sum_{i=1}^{n_{y}}{\left(\overline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k}\left(i\right)\right)}$
$=$ $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right),$
where $\left(\overline{Y}_{k|k-1}\left(i\right)-\overline{Y}_{k|k}\left(i\right)+\underline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)$ is the length of i-th interval of $\Delta\mathcal{Y}_{k|k}$.
According to Proposition 4, to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, it is equivalent to reduce the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ since the amount of prior uncertainty $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$ is fixed at time $k$. As shown in the next theorem, the amount of uncertainty reduction, i.e., $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, is bounded by the new information extracted from $\mathcal{M}^{x}_{k|k}$.
Theorem 5.
The amount of uncertainty reduction at $k$ is
$\!\!\!\!\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}+\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (38)
satisfying
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\geqslant 2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1},$
$\!\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant\!\left\|\!A_{3}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
where $\Delta\mathcal{X}_{k-1|k}=\mathcal{X}_{k-1|k-1}-\mathcal{X}_{k-1|k}$ and $\Delta\mathcal{Y}_{k-1|k}=\mathcal{Y}_{k-1|k-1}-\mathcal{Y}_{k-1|k}$ are the adversary’s uncertainty reduction of $X_{k-1}$ and $Y_{k-1}$, and $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$ quantifies the difference in central estimation with and without considering $\mathcal{X}_{k|k}$, $\left\|A\right\|\frac{\Delta}{=}\sum_{i,j}^{n,m}{\left|a_{i,j}\right|}$ is the absolute value of $A$.
Proof 3.4.
See Appendix C.
According to Theorem 5, the reduction of uncertainty $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is highly correlated with the amount of information that the adversary extracts from the observation set $\mathcal{M}_{k|k}^{x}$. Furthermore, with Proposition 4 and Theorem 5, we have the following lemma to bound the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k-1|k}\right)$.
Lemma 6.
The privacy level can be bounded with the following inequalities,
$\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
$\leqslant\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k}\!\right)\!\leqslant\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$
Therefore, on the one hand, one can reduce the extracted information $\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$. On the other hand, when the private level is high, we also avoid the adversary to update its central estimation $c_{k|k}^{y}$ due to the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$.
4 Privacy Filter Design Problem Using the Volumetric Privacy measure
In this section, we address the privacy filter design problem, where the filter outputs an appropriate observation set that balances the trade-off between the data utility of the public state and the privacy protection of the private state.
4.1 The Structure of Privacy Filter
We begin by defining the decision domain of the privacy filter as follows. At time $k$, given the last decision set $\mathcal{X}_{k-1|k-1}$ and the private set $\mathcal{Y}_{k-1|k-1}$, the tightest interval of $X_{k}$ can be computed via Lemma 9, i.e.,
$\!\!\mathcal{X}_{k|k-1}\!=\!\Psi\!\left(\!A_{1}\!\right)\!\mathcal{X}_{k-1|k-1}\!+\!\Psi\!\left(\!A_{2}\!\right)\!\mathcal{Y}_{k-1|k-1}\!+\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x}\!,$
which contains all possible public states that can be reached from any states in $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$. Therefore, $\mathcal{X}_{k|k-1}$ is the maximum observation set $\mathcal{M}^{x}_{k|k}$ that the filter can release, i.e., $\mathcal{M}^{x}_{k|k}\subset\mathcal{X}_{k|k-1}$. To maintain the high data utility, the observation set has to satisfy the following constraint,
$\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}.$
We next consider the privacy filter drawn in Fig.2, which first randomly generates a set $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$ that contains the actual public state $X_{k}=x_{k}$, and then optimizes the observation set $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ based on linear programming. We will demonstrate that the adversary cannot reduce its uncertainty of the public state by reversing the proposed filtering policy.
Figure 2: The structure of privacy filter.
4.2 Randomization
We consider the following random set
${\mathcal{S}}_{k|k}^{x}=\left[\begin{array}[]{c}x_{k}-{\alpha}_{k}\left(x_{k}-\underline{X}_{k|k-1}\right)\\ x_{k}+{\beta}_{k}\left(\overline{X}_{k|k-1}-x_{k}\right)\\ \end{array}\right],$ (41)
where $\alpha_{k}$ and $\beta_{k}$ are uniform random variables with
${\alpha}_{k}\in\left[0,\frac{\epsilon^{x}}{2\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}}\right],$
and
${\beta}_{k}\in\left[0,\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\right].$
Since $\left(x_{k}-\underline{X}_{k|k-1}\right)$ is the radius between the actual public state $X_{k}=x_{k}$ and the lower endpoint of $\mathcal{X}_{k|k-1}$, and $\left(\overline{X}_{k|k-1}-x_{k}\right)$ is the radius between $x_{k}$ and the upper endpoint of $\mathcal{X}_{k|k-1}$, the random set ${\mathcal{S}}_{k|k}^{x}$ is a subset of $\mathcal{X}_{k|k-1}$ that contains the actual public state. Also, we have $\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ as
$\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)$
$=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
$\leqslant$ $\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}$
$+\frac{\epsilon^{x}}{2\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
$=$ $\epsilon^{x}.$
We next restrict $\mathcal{S}^{x}_{k|k}$ be the subset of the observation set $\mathcal{M}^{x}_{k|k}$, and optimize $\mathcal{M}^{x}_{k|k}$ to enhance the privacy level.
4.3 Privacy Filter Optimization
[…]
$\mathbf{P_{1}}:$ $\max_{\mathcal{M}^{x}_{k|k}}\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ (42)
[…]
where $\mathcal{S}^{x}_{k|k}\subseteq\mathcal{M}^{x}_{k|k}$ is required by the structure of the privacy filter, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ describes the prior space of $\mathcal{M}^{x}_{k|k}$, and $\mathrm{Vol}\left({\mathcal{M}}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ enforces the requirement of data utility. We next demonstrate that the optimization problem $\mathbf{P_{1}}$ can be solved via linear programming.
Theorem 7.
The privacy filter optimization problem $\mathbf{P_{1}}$ is equivalent to the following linear programming
$\mathbf{P_{2}}:\max_{\epsilon^{y},\mathcal{M}^{x}_{k|k},p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}}\epsilon^{y}$ (48)
[…]
$\left\{\begin{array}[]{c}p_{k-1|k}^{\Delta z}\geqslant 0\\ p_{k-1|k}^{\Delta z}\geqslant p_{k-1|k-1}^{z}-p_{k-1|k}^{m,z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\\ \end{array}\right.,\text{for}\;Z=X,Y$ (59)
where $\epsilon^{y}\geqslant 0$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{R}^{2n_{x}}$ and $p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}\in\mathcal{R}^{n_{x}}$.
Proof 4.1.
See Appendix D
Figure 3: Inference attack after the optimal privacy filter with $\mathrm{Vol}\left(\mathcal{M}_{k|k}^{x}\right)\leqslant 0.01$.
Figure 4: Inference attack after the optimal privacy filter with $\mathrm{Vol}\left(\mathcal{M}_{k|k}^{x}\right)\leqslant 0.5$.
As a result, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 2.3. Although the linear programming problem $\mathbf{P_{2}}$ is deterministic, we can demonstrate that the adversary cannot reduce its uncertainty by reversing the optimization process.
Proposition 8.
If the adversary infers the private state via the inference attack (2)-(32), then it cannot find a smaller interval subset of $\mathcal{M}_{k|k}^{x,\star}$ containing the actual public state $X_{k}=x_{k}$ via reverting the linear programming.
Proof 4.2.
On the one hand, the actual public state $X_{k}=x_{k}$ can be any element of $\mathcal{X}_{k|k-1}$, and $S_{k}^{x}$ is a random subset of $\mathcal{X}_{k|k-1}$ containing $x_{k}$. Thus, the upper or lower endpoint of $S_{k}^{x}$ may coincide with $x_{k}$.
On the other hand, since $\mathcal{M}_{k|k}^{x,\star}$ is the optimal observation set obtained from the linear programming, for any $S_{k}^{x}$ in the vicinity of $\mathcal{M}_{k|k}^{x,\star}$, i.e., $S_{k}^{x}\in\left\{S_{k}^{x}\left|\lim_{\left\|\Delta\right\|_{1}\rightarrow 0}{S_{k}^{x}+\Delta=\mathcal{M}_{k|k}^{x,\star}}\right.\right\}$, the optimal solution remains $\mathcal{M}_{k|k}^{x,\star}$.
As a result, it is possible that the upper or lower endpoint of $\mathcal{M}_{k|k}^{x,\star}$ corresponds to the actual public state. Thus, the minimal interval observation set for the adversary is $\mathcal{M}_{k|k}^{x,\star}$.
According to Proposition 8, even if the adversary knows the structure of the privacy filter, it cannot revert the linear programming to obtain a smaller interval containing $x_{k}$ to reduce uncertainty. In other words, the privacy filer is robust against the reverse optimization attack.
5 Numerical Verification
In this section, we study the performance of privacy filter for the production-inventory problem with the following parameters
$A_{1}=\left[\begin{matrix}1.00&0.00\\ 0.00&1.00\\ \end{matrix}\right],A_{2}=\left[\begin{matrix}0.40&0.80\\ 0.60&0.20\\ \end{matrix}\right],$
$A_{3}=\left[\begin{matrix}0.50&-0.90\\ -0.10&-0.10\\ \end{matrix}\right],A_{4}=\left[\begin{matrix}-0.10&-0.90\\ 0.10&0.00\\ \end{matrix}\right],$
$B_{1}=\left[\begin{matrix}-1.00&0.00\\ 0.00&-1.00\\ \end{matrix}\right],B_{2}=\left[\begin{matrix}4.20&0.00\\ 0.00&2.40\\ \end{matrix}\right],$
$\left(\mathcal{W}_{k}^{x}\right)^{\top}=\left[\begin{matrix}1.74&1.91&1.94&2.01\\ \end{matrix}\right],$
$\left(\mathcal{W}_{k}^{y}\right)^{\top}=\left[\begin{matrix}0.91&0.23&0.95&0.43\\ \end{matrix}\right].$
The initial state sets are assumed to be
$\left(\mathcal{X}_{0|-1}\right)^{\top}=\left[\begin{matrix}1.00&0.24&1.20&0.40\\ \end{matrix}\right],$ (60)
$\left(\mathcal{Y}_{0|-1}\right)^{\top}=\left[\begin{matrix}2.40&0.60&3.70&1.30\\ \end{matrix}\right].$ (61)
In our simulation, the initial states are uniformly sampled from the bounded sets (60)-(61). To simulate the approximate periodic fluctuations in demand and productivity, the actual process noises are set to be
$\left(W_{k}^{x}\right)^{\top}=\left[\begin{matrix}1.88+0.03\cos\left(\frac{2\pi k}{30+7\rho_{k}}\right)&1.94\\ \end{matrix}\right],$
$\!\!\left(W_{k}^{y}\right)\!^{\top}\!=\!\left[\!\begin{matrix}0.944\!+\!0.006\cos\left(\!\frac{2\pi k}{7+2\gamma_{k}}\!\right)\!\!&\!\!0.33+0.094\sin\left(\!\frac{2\pi k}{7+4\tau_{k}}\!\right)\\ \end{matrix}\!\right]\!,$
where $\rho_{k}$, $\gamma_{k}$ and $\tau_{k}$ are uniform random variables in $\left[0,1\right]$. As discussed in Section 2, the production rate is private but the inventory information has to be released.
We first plot the trajectories of system states and their interval tubes in Fig.3 and Fig.4 under the optimal privacy filter design for different values of $\epsilon^{x}$. We also use the central point of the posterior intervals as one of the possible testing estimation, e.g., $\frac{\overline{X}_{k|k}+\underline{X}_{k|k}}{2}$ for $\mathcal{X}_{k|k}$. The pink areas in these figures represent the uncertainty sets of system states. As shown in Fig.3, when $\epsilon^{x}=0.01$ is small, the adversary’s uncertainty about the private production rate is small, and its central estimation closely matches the actual production rate. However, when $\epsilon^{x}$ increases to $0.5$, the utility of the inventory information decreases slightly, but this leads to higher uncertainty of the inference attack, causing the adversary’s estimation of the production rate to become less accurate. Therefore, the proposed privacy filter effectively reduces the privacy leakage of the production rate, though at the cost of introducing some inaccuracies in the inventory information.
Figure 5: The truncated Gaussian mechanism.
We analyze the utility-privacy trade-off performance of the optimal privacy filtering policy and compare it with the truncated Gaussian mechanism, which is proposed in [18] for differential privacy. In the truncated Gaussian mechanism, illustrated in Fig. 5, the original state $x_{k}$ is perturbed by additive noise $v_{k}$ drawn from a truncated Gaussian distribution defined over the interval $\left[-\epsilon^{x}/2,\epsilon^{x}/2\right]$, with zero mean and variance $\left(\epsilon^{x}\right)^{2}$. Subsequently, the perturbed observation set is released as follows:
$\mathcal{M}_{k|k}^{x}=\left[\begin{array}[]{c}z_{k}-\epsilon^{x}/\left(2n_{x}\right)\cdot\mathbf{1}_{n_{x}}\\ z_{k}+\epsilon^{x}/\left(2n_{x}\right)\cdot\mathbf{1}_{n_{x}}\\ \end{array}\right].$ (62)
Since the noise $v_{k}$ is constrained within $\left[-\epsilon^{x}/2,\epsilon^{x}/2\right]$, the released public state $x_{k}$ is guaranteed to reside within the interval $\mathcal{M}_{k|k}^{x}$, and the utility of $x_{k}$ is quantified by $\epsilon^{x}$.
The privacy-utility trade-off is evaluated by plotting the privacy level of the production rate against the utility of the inventory under different approaches, as depicted in Fig. 6. The results indicate that increasing the utility of inventory leads to a reduction in the privacy level of the production rate, highlighting the inherent trade-off between data utility and privacy protection. Compared to the truncated Gaussian mechanism, the optimal privacy filter achieves a higher privacy level while maintaining lower data distortion.
Figure 6: The privacy level of the private state and utility of the public state.
6 Conclusion
In this paper, we developed a volumetric privacy measure for dynamic systems with UBB noise. We defined the inference attack that an adversary uses to estimate the private state and introduced a volumetric measure to evaluate the privacy level. We provided computational approaches based on interval analysis and analyzed the theoretical properties of the proposed measure. Furthermore, we proposed an optimization-based approach for privacy filter design to defend the system against inference attacks. The effectiveness of our method was demonstrated through a production-inventory case study.
Appendix A Proof of Lemma 2
At the time step $k=0$, the adversary only has prior knowledge, i.e., $\mathcal{Y}_{0|-1}$, therefore, its inference set is $\mathcal{Y}_{0|0}=\mathcal{Y}_{0|-1}$.
To prove Lemma 2 for $k\geqslant 1$, we need the following lemma that computes the tightest interval by forward reachability analysis.
Lemma 9.
[1, 2] Consider the static system $S=AM+BW$, where $M$ and $W$ are bounded intervals, the tightest interval for $S$, i.e., $\mathcal{S}$ can be computed as
$\mathcal{S}=\Psi\left(A\right)\mathcal{M}+\Psi\left(B\right)\mathcal{W},$ (63)
where
$\Psi\left(N\right)=\left[\begin{matrix}\frac{N+\left|N\right|}{2}&\frac{N-\left|N\right|}{2}\\ \frac{N-\left|N\right|}{2}&\frac{N+\left|N\right|}{2}\\ \end{matrix}\right].$ (64)
Also, we can compute the radius and center of the intervl $\mathcal{S}$ via
$p^{s}=\left|A\right|p^{m}+\left|B\right|p^{w},$ (65)
$c^{s}=Ac^{m}+Bc^{w}.$ (66)
When $A_{1}$ and $A_{2}$ are invertible, we have
$X_{k-1}=A_{1}^{-1}X_{k}+\left(-A_{1}^{-1}A_{2}\right)Y_{k-1}+\left(-A_{1}^{-1}B_{1}\right)W_{k}^{x},$
and
$Y_{k-1}=A_{2}^{-1}X_{k}+\left(-A_{2}^{-1}A_{1}\right)X_{k-1}+\left(-A_{2}^{-1}B_{1}\right)W_{k}^{x}.$
According to Lemma 9, the tightest intervals for (2.3) and (2.3) are (2) and (2). Then, the intersection of different intervals, i.e., (12) and (13), can be computed with (24) and (27). Finally, the one-step forward reachable set (14) can be approximated with the tightest interval (32) based on Lemma 9, and the calibrated uncertainty set $\mathcal{X}_{k|k}$ and the tightest prior inference set $\mathcal{Y}_{k|k-1}$ can be approximated similarly. Finally, since at the time step $k=0$, the backward calibration (2.3) and (2.3) is not available, the adversary can only calibrate the public state set with its prior knowledge $\mathcal{X}_{0|-1}$ and the observation set $\mathcal{M}_{0|0}^{x}$ according to (36).
Appendix B Proof of Lemma 3
According to Lemma 9, the radius of $\mathcal{M}^{y}_{k-1|k}$ can be computed as
$p_{k-1|k}^{m,y}=\left|A_{2}^{-1}\right|p_{k|k}^{m,x}+\left|A_{2}^{-1}A_{1}\right|p_{k-1|k-1}^{x}+\left|A_{2}^{-1}B_{1}\right|p_{k}^{w,x},$
where $p^{m,x}_{k|k}$, $p_{k-1|k-1}^{x}$ and $p^{w,x}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{X}_{k-1|k-1}$ and $\mathcal{W}_{k}^{x}$, respectively. Since $\mathcal{Y}_{k-1|k}$ is the intersection result from $\mathcal{M}^{y}_{k-1|k}$ and $\mathcal{Y}_{k-1|k-1}$, the radius of $\mathcal{Y}_{k-1|k}$ is smaller than the radius of $\mathcal{M}^{y}_{k-1|k}$, i.e., $p_{k-1|k}^{y}\leqslant p_{k-1|k}^{m,y}$. Also, the radius of $\mathcal{Y}_{k|k}$ can be computed as
$p_{k|k}^{y}=\left|A_{3}\right|p_{k-1|k}^{x}+\left|A_{4}\right|p_{k-1|k}^{y}+\left|B_{2}\right|p_{k}^{w,y}.$
Therefore, with (B), we have
$p_{k|k}^{y}\leqslant\left|A_{3}\right|p_{k-1|k}^{x}+\left|B_{2}\right|p_{k}^{w,y}+\left|A_{4}\right|\left|A_{2}^{-1}B_{1}\right|p_{k}^{w,x}$
$+\left|A_{4}\right|\left(\left|A_{2}^{-1}\right|p_{k|k}^{m,x}+\left|A_{2}^{-1}A_{1}\right|p_{k-1|k-1}^{x}\right).$
Since $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$ and $\mathcal{X}_{k-1|k}$ is a subset of $\mathcal{M}_{k-1|k-1}^{x}$, we have $p_{k-1|k}^{x}\leqslant p^{m,x}_{k-1|k-1}\leqslant\overline{p}^{x}$ for any $k\geqslant 1$, thus we have (3).
Appendix C Proof of Theorem 5
The difference set $\Delta\mathcal{Y}_{k|k}$ is computed as,
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}=\Phi\left(A_{3}\right)\Delta\mathcal{X}_{k-1|k}+\Phi\left(A_{4}\right)\Delta\mathcal{Y}_{k-1|k},$
where
$\Delta\mathcal{X}_{k-1|k}$ $=\mathcal{X}_{k-1|k-1}-\mathcal{M}_{k-1|k}^{x}$
$=\left[\begin{array}[]{c}\min\left\{\underline{X}_{k-1|k-1}-\underline{M}_{k-1|k}^{x},0\right\}\\ \max\left\{\overline{X}_{k-1|k-1}-\overline{M}_{k-1|k}^{x},0\right\}\\ \end{array}\right],$ (69)
$\Delta\mathcal{Y}_{k-1|k}$ $=\mathcal{Y}_{k-1|k-1}-\mathcal{Y}_{k-1|k}$
$=\left[\begin{array}[]{c}\min\left\{\underline{Y}_{k-1|k-1}-\underline{M}_{k-1|k}^{y},0\right\}\\ \max\left\{\overline{Y}_{k-1|k-1}-\overline{M}_{k-1|k}^{y},0\right\}\\ \end{array}\right].$ (72)
Therefore, the volume of the difference set is (38).
With Lemma 9, we have
$p_{k|k}^{\Delta y}=\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y},$ (73)
where the radius $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ can be computed via
$2p_{k-1|k}^{\Delta z}$
$=$ $\max\left\{\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},0\right\}-\min\left\{\underline{Z}_{k-1|k-1}-\underline{M}_{k-1|k}^{z},0\right\}$
$=$ $\max\left\{\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},0\right\}+\max\left\{\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1},0\right\}$
$=$ $\max\left\{0,\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}+\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1},\right.$
$\left.\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\right\},\text{for}\,Z=X,Y,$ (74)
which satisfies $p_{k-1|k}^{\Delta z}\geqslant 0$. As a result, we have
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$
$=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
=$\overset{(a)}{=}\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
$\leqslant\left\|A_{3}\right\|\left\|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|A_{4}\right\|\left\|p_{k-1|k}^{\Delta y}\right\|_{1}$
$=\left\|A_{3}\right\|\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right),$ (75)
where $(a)$ is due to $p_{k-1|k}^{\Delta x}\geqslant 0$ and $p_{k-1|k}^{\Delta y}\geqslant 0$.
Besides, given an interval $\mathcal{X}$, we can express it with its center point and radius, i.e., $\mathcal{X}=\left[\begin{array}[]{c}\frac{c-p}{2}\\ \frac{c+p}{2}\end{array}\right]$. Therefore, we have (C),
$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$
$=$ $\left\|\overline{Y}_{k|k-1}-\overline{Y}_{k|k}\right\|_{1}+\left\|\underline{Y}_{k|k-1}-\underline{Y}_{k|k}\right\|_{1}$
$=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
$\geqslant$ $\left\|\overline{Y}_{k|k}+\underline{Y}_{k|k}-\left(\overline{Y}_{k|k-1}+\underline{Y}_{k|k-1}\right)\right\|_{1}$
$\geqslant$ $2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$ (76)
Appendix D Proof of Theorem 7
To maximize the privacy level, it is equivalent to minimize the amount of uncertainty reduction since we have $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, where the prior uncertainty set $\mathcal{Y}_{k|k-1}$ is fixed at time step $k$.
Besides, the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=\left\|p_{k|k}^{\Delta y}\right\|_{1}=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$, where the elements of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ are non-negative vectors as shown in (C). Therefore, we can replace the objective function with the slack variable $\epsilon^{y}$ and add $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\leqslant\epsilon^{y}$ as a new constraint, and then minimize $\epsilon^{y}$.
Since $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is determined by $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, we can replace constraints (24) and (27) with the constraints of difference sets (C). Also, the objective function increases with any elements of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ since the elements of $p_{k-1|k}^{\Delta x}$, $p_{k-1|k}^{\Delta y}$, $\left|A_{3}\right|$ and $\left|A_{4}\right|$ are non-negative. As a result, we can replace the constraint of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, i.e., (C), with inequalities (59), and let $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ be decision variables.
Besides, the constraints $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ and $\overline{M}^{x}_{k|k}\geqslant\underline{M}^{x}_{k|k}$ are equivalent to the inequality constraint, $\underline{X}_{k|k-1}\leqslant\underline{M}^{x}_{k|k}\leqslant\underline{S}^{x}_{k|k}\leqslant\overline{S}^{x}_{k|k}\leqslant\overline{M}^{x}_{k|k}\leqslant\overline{X}_{k|k-1}$, and the utility constraint $\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ can be replaced with $\left\|\overline{M}^{x}_{k|k}-\underline{M}^{x}_{k|k}\right\|_{1}\leqslant\epsilon^{x}$.
Finally, the objective and the constraints are linear functions of the decision variables, thus, the optimal privacy filter can be obtained by solving the linear programming $\mathbf{P_{2}}$.

### v4

@@
-A Volumetric Approach to Privacy of Dynamical Systems
+A Volumetric Privacy Measure for Dynamical Systems With Bounded Disturbance
 Abstract
-Information-theoretic metrics, such as mutual information, have been widely used to evaluate privacy leakage in dynamic systems.
-However, these approaches are typically limited to stochastic systems and face computational challenges.
-In this paper, we introduce a novel volumetric framework for analyzing privacy in systems affected by unknown but bounded noise.
-Our model considers a dynamic system comprising public and private states, where an observation set of the public state is released.
-An adversary utilizes the observed public state to infer an uncertainty set of the private state, referred to as the inference attack.
-We define the evolution dynamics of these inference attacks and quantify the privacy level of the private state using the volume of its uncertainty sets.
-We then develop an approximate computation method leveraging interval analysis to compute the private state set.
-We investigate the properties of the proposed volumetric privacy measure and demonstrate that it is bounded by the information gain derived from the observation set.
-Furthermore, we propose an optimization approach to designing privacy filter using randomization and linear programming based on the proposed privacy measure.
-The effectiveness of the optimal privacy filter design is evaluated through a production-inventory case study, illustrating its robustness against inference attacks and its superiority compared to a truncated Gaussian mechanism.
+This paper presents a volumetric privacy framework for dynamical systems subject to bounded disturbances, developed without requiring prior knowledge of their probability distributions.
+We consider systems with both public and private states, where a set containing the public state is shared as the observation.
+An adversary is assumed to execute an inference attack by exploiting the observed public state set to estimate an uncertainty set for the private state.
+The volume of this inferred set quantifies the adversary’s estimation uncertainty and serves as the proposed volumetric privacy metric.
+Approximate set-membership estimation techniques are developed to compute the private-state uncertainty set, and the properties of the privacy measure are analyzed, demonstrating that it is bounded by the information gain from the observation set.
+Furthermore, an optimization-based privacy filter design problem is formulated, employing randomization and linear programming to enhance the volumetric privacy level.
+The effectiveness of the proposed approach is validated through a production–inventory case study.
+Results show that the optimal privacy filter significantly improves robustness against inference attacks and outperforms two baseline mechanisms based on additive noise and quantization.
 keywords:
-Volumetric privacy measure; privacy protection; interval analysis; truncated Gaussian mechanism.
+Volumetric privacy measure; privacy protection; interval analysis; bounded disturbance.
 thanks: The work was partially supported by the Research Grants Council of Hong Kong under Project CityU 21208921, a grant from Chow Sang Sang Group Research Fund sponsored by Chow Sang Sang Holdings International Limited.
@@
 However, the exposure of such data may inadvertently reveal sensitive information.
-Specifically, correlations between shared metrics and underlying operational parameters can be exploited by adversaries to develop competitive and malicious strageties.
-This challenge highlights the critical need for methodologies that preserve data utility while ensuring rigorous privacy protection for dynamic systems.
-Information-theoretic metrics, including mutual information, directed information, and conditional entropy, have been extended to quantify privacy leakage in dynamic systems.
-Despite their utility, these metrics have notable limitations.
-For instance, as highlighted in [8, 30], mutual information evaluates privacy leakage in an average sense, which can overlook privacy breaches for infrequent realizations of private variables.
-The computational complexity of these measures is another challenge, as they require evaluating leakage for all possible realizations of random variables [25, 32, 11].
-Most existing information-theoretic approaches are focused on stochastic systems and typically assume complete knowledge of the probability distributions of system states.
-However, this assumption does not hold in systems that are influenced by unknown but bounded (UBB) noise.
-To address these challenges, this paper introduces a volumetric privacy measure based on the uncertainty set of the private variable.
-The proposed approach is applicable to both deterministic and stochastic systems and eliminates the need for prior knowledge of probability distributions.
+Specifically, correlations between shared metrics and underlying operational information can be exploited by adversaries to develop competitive and malicious strageties.
+This challenge highlights the critical need for methodologies that preserve data utility while ensuring privacy protection for dynamic systems.
+Dynamical systems subject to bounded disturbances without knowledge of their underlying distributions provide a natural framework for modeling numerous practical applications involving sensitive information, e.g., inventory–production systems and connected vehicles.
+However, the notion of privacy in such systems remains insufficiently explored.
+Within the context of set-membership estimation, the states of these systems can be represented by geometric sets, such as ellipsoids or zonotopes, whose volumes quantify the degree of inference uncertainty.
+Motivated by these considerations, this paper investigates the notion of volumetric privacy for systems affected by bounded disturbance.
+We develop privacy-preserving strategies aimed at maximizing an adversary’s uncertainty, i.e., the volume of uncertainty set, in inferring private states.
+The proposed approaches are applicable to both deterministic and stochastic systems, without requiring prior knowledge of the underlying probability distributions.
+1.2 Related Work
+Stochastic approaches to privacy primarily include differential privacy and information-theoretic methods.
+Differential privacy (DP) [1] has been incorporated into dynamic settings through differentially private Kalman filtering [2], DP-preserving average consensus via noise injection [3], and minimal-noise mechanisms for multi-agent systems based on observability properties [4].
+Recent work [5] introduced a trace-based variance–expectation ratio to quantify topology preservation and derived optimal noise designs, while [6] provided a comprehensive survey of DP in dynamical systems.
+In parallel, information-theoretic approaches quantify privacy leakage using mutual information or conditional entropy.
+Recent studies include mutual-information-based private filtering for hidden Markov models [7], directed-information-based privacy filters for linear systems [8, 9], and recent extensions to partially observable Markov decision processes addressing privacy-aware estimation and control [10, 11].
+Most existing studies focus on deterministic or stochastic systems with unbounded noise and known distributions, leaving privacy protection for systems subject to unknown-but-bounded disturbances relatively unexplored.
+Recent works [12, 13] developed differentially private set-based estimators using truncated noise, while [14] introduced guaranteed privacy concepts and optimization methods for $\mathcal{H}_{\infty}$-based privacy-preserving interval observers.
+State-opacity-based methods [15, 16] ensured indistinguishable outputs between secret and non-secret states but did not quantify the associated estimation uncertainty.
+Note that set-membership estimators typically characterize uncertainty through bounded geometric sets such as intervals [17], zonotopes [18], or ellipsoids [19], where the corresponding set volume naturally describes the amount of estimation uncertainty.
+Motivated by this, we analyze privacy leakage in systems with private and public states and propose a volumetric approach that maximizes the private-state set volume, thereby enhancing privacy while explicitly accounting for geometric effects under inference attacks.
+There are some related deterministic approaches to privacy without adding noise, e.g., plausible deniability [20] and noiseless privacy [21].
+In [20], privacy leakage in deterministic systems was measured by the volume of reachable state sets, and was determined by the observability.
+However, this framework does not apply to systems with bounded disturbance, where inference uncertainty depends on both observability and disturbance.
+Moreover, the problem of privacy filter design was not addressed in [20], whereas we propose a concrete design using randomization and optimization.
+The work in [22] addressed parameter privacy in deterministic systems via constrained convex generators (CCGs), which differs from our private state protection setting.
+Also, defense strategies in [22] involve ceasing information sharing or altering parameters, which might be unsuitable for fixed-parameter systems with continuous communication.
+As discussed in Sec.
+3.1, the complexity of CCG-based inference grows exponentially with time, motivating our use of interval analysis for computational efficiency.
+Noiseless privacy [21] and non-stochastic privacy [23] employed non-stochastic information-theoretic approaches to limit information leakage, assuming static private-variable domains and without accounting for temporal dependencies in sequential data.
+While noiseless privacy, non-stochastic privacy, and our volumetric privacy all achieve privacy through the release of bounded outputs.
+However, in our setup, dynamical systems subject to bounded disturbance have time-variant private-state reachable sets that can be recursively estimated, enabling dynamic leakage evaluation.
+Building on this insight, the proposed volumetric privacy filter dynamically evaluates the private state set and adapts the observation accordingly, thereby achieving higher privacy levels with lower data distortion, as shown in Sec.
+5.
+Finally, other deterministic privacy-preserving approaches often exploit observability reduction or state decomposition to protect private information in multi-agent systems.
+For example, the authors in [24] established a connection between network privacy and its observability space, proposing a privacy-aware communication protocol that achieves average consensus while protecting initial states.
+In [3], the trace of the observability Gramian was employed to quantify information leakage through an intruder node, and an online optimization strategy was proposed to adapt communications in order to minimize such leakage.
+Privacy-preserving consensus designs were also proposed in [25, 26], including augmented states and novel consensus algorithms with simultaneous accuracy, resilience, and privacy guarantees.
+State decomposition methods, as in [27], split each node’s state into randomized components to prevent disclosure of individual states during consensus.
 1.3 Contributions
-This paper addresses the privacy protection problem for dynamic systems with UBB noise, as illustrated in Fig.
-1.
-In this framework, the system state is partitioned into two categories: the public state $X_{k}$ and the private state $Y_{k}$, both of which belong to a known bounded set.
-The observation of the system states is represented by the set $\mathcal{M}_{k|k}^{x}$ which contains the actual public state $X_{k}=x_{k}$.
-An untrusted third party, i.e., the adversary, uses $\mathcal{M}_{k|k}^{x}$ to infer the private state by constructing an uncertainty set $\mathcal{Y}_{k|k}$, which is referred to as the inference attack.
-Figure 1: The inference attack.
-This system setup is inspired by the production-inventory problem discussed in Section 2.
-In such scenarios, inventory information can aid distributors in optimizing sales.
-However, it also enables adversaries to infer sensitive information, such as the production rate, which reveals a company’s private data, including production efficiency, strategies, and supply chain dynamics.
-The contributions of this paper are summarized as follows:
-  • Definition of Inference Attack: We define the adversary’s inference attack using set operations.
-  • Volumetric Privacy Measure: A novel privacy measure is proposed based on the volume of the uncertainty set for the private state.
-  • Privacy Level Computation: We develop a computational method utilizing interval analysis to evaluate the privacy level.
-  • Optimal Privacy Filter Design: An optimization-based approach is introduced to design privacy filters, based on randomization and linear programming.
-Additionally, we validate our approach using the production-inventory example.
-The results demonstrate that the adversary can infer sensitive production rates, through inventory data.
-However, this leakage can be significantly mitigated under the proposed optimal privacy filter design.
-Also, compared with the truncated Gaussian mechanism, our privacy filter design achieves lower data distortion while maintaining the same privacy level.
-1.3 Related work
-Differential privacy was initially defined to hide the presence of an individual or a specific record in the responses of queries for static datasets.
-The authors in [15] extended the differential privacy to dynamic systems, and proposed the differentially private Kalman filter for linear systems.
-In [5], the authors proposed a general architecture for private filtering by approximating the desire filter in a differential private way, and studied the application of differential privacy in distributed control and optimization.
-In [21], the authors studied the differential privacy preserving average consensus problem to protect the initial state in distributed control by adding and abstracting random noises.
-The authors in [12] connected the input observability with differential privacy and proposed the privacy preserving controller design method.
-The authors in [33] studied the minimal amount noises added to multi-agent systems for differential privacy based on the minimal observability subspace.
-They proposed node-based and edge-based privacy-preserving mechanisms based on the optimization of the minimum added noise, and discussed the trade-offs between utility and privacy.
-In [34], the authors proposed several cost-friendly differential privacy-preserving schemes by randomly adjusting the charge-discharge rate of a batter based on a modified Laplace distribution.
-Information-theoretic protection methods use conditional entropy, mutual information, etc, from information theory, to measure privacy leakage.
-In [22], the authors studied the privacy filter design for a Markov chain to hide the private state when transmitting the monitored state to a receiver, which was approximately solved with a greedy algorithm based on the convex optimization.
-The authors in [3] designed a private filter to protect the state of a hidden Markov chain via minimizing the mutual information with the constraint of the observation utility.
-In [31, 24], the authors studied the privacy-aware filter design for linear systems, using the directed information between the system state and the output of the filter as the privacy measure.
-It was proved that the optimal filter can be realized with a Kalman filter and an additive Gaussian mechanism.
-The authors in [32] studied the structural properties of the privacy-aware state estimation problem based on the dynamic programming decomposition.
-They approximated the privacy leakage with variational techniques and solved the estimation problem via policy gradient approaches.
-In [23], the conditional entropy of the private state given measurements and controls was used as a penalty term in the partially observable Markov decision process (POMDP) optimization, to avoid the adversary accurately estimate the system state.
-They showed that the state obfuscation problem is a standard POMDP optimization problem in which the cost-to-go function is concave in the belief state.
-In [16], the authors considered to protect the privacy of the electricity demand of a household to avoid the adversary infers the indoor activity.
-They utilized the rechargeable battery to hide the actual electricity demand and obtained the optimal charging policy via minimizing the privacy leakage measured by mutual information.
-However, the classical information-theoretic metrics for privacy might have some limitations when facing one-try attack as discussed in [30], which motivates the development of $R\acute{e}nyi$ min-entropy [30], maximal leakage [9] and pointwise maximal leakage [28], etc.
-In most existing privacy protection literature for dynamic systems, it is assumed that the system are disturbed with unbounded noise that has probabilistic properties, e.g., Gaussian distribution.
-However, there are also dynamic systems are driven with UBB noise without probability assumptions.
-The privacy concept for these systems are not well developed.
-The authors in [6, 7] designed differentially private set-based estimator to protect the privacy of measurements based on a truncated noise distribution.
-In [13], the authors proposed the guaranteed privacy extent from differential privacy to hide the measurement sensor identities, and provided the optimization method for the $\mathcal{H}_{\infty}$ optimal privacy preserving interval observer design.
-Furthermore, in [26, 27], authors studied the state opacity problems where the discrete state space is divided into disjoint sets of secrete and non-secrete states, and the system outputs from secrete and non-secrete sets are indistinguishable.
-The authors in [19] proposed $\delta-$approximate initial-state opacity as a secrete notion for discrete-time systems with continuous states, and developed verification approaches based on Barrier certificates.
-The proposed volumetric privacy measure in this work is inspired by set-membership estimation methods for systems with UBB noise [20].
-Most set-membership methods represent states as bounded geometric sets, such as intervals [10], zonotopes [14], or ellipses [4], where the volumes of these sets correspond to estimation uncertainty.
-Building on these methods, we consider the adversary’s ability to infer the private state set based on observations of the public state.
-Consequently, a larger volume of the uncertainty set implies a broader range of potential private states that could be consistent with the observed public state set.
-This motivates our use of the volume of the private state set as a privacy measure.
-This work differs from prior studies, such as [6], [7], and [13], which focus on ensuring the differential privacy of system states.
-In contrast, we address the problem of protecting private states while publicly releasing non-sensitive states, as illustrated in Fig.
-1.
-Our approach uses a volumetric privacy measure, which also distinguishes it from state opacity problems [26], [27], and [19], where the state space is partitioned into secret and non-secret sets.
-Furthermore, state opacity problems focus on guaranteeing privacy by ensuring that at least one non-secret state produces an observation indistinguishable from the secret state.
-In these frameworks, enlarging the volume of the non-secret state set does not enhance privacy, as privacy is determined solely by the existence of indistinguishable outputs, not by the size of the state set.
-Differently, within our framework, a larger private state set is preferred, as it increases the adversary’s uncertainty in inferring the private state.
+This paper studies privacy for dynamic systems with bounded disturbance as shown in Fig.
+1, where the system state is split into public $X_{k}$ and private $Y_{k}$, and an adversary uses the public observation set $\mathcal{M}_{k|k}^{x}$ to infer $Y_{k}$ via an uncertainty set $\mathcal{Y}_{k|k}$.
+Figure 1: The system setup.
+The primary contribution of this work is the development of an extensible framework for privacy analysis and mitigation in dynamic systems subject to bounded disturbance.
+This contribution can be summarized in three principal aspects.
+(1) Volumetric Privacy Measure: We introduce a privacy metric based on the volume of the estimated private-state set obtained via set-membership estimation given the available observations.
+(2) Privacy Level Computation: We develop computational methods to quantify privacy level and prove that the relevant privacy leakage is bounded by the information gain from the observations.
+(3) Optimal Privacy Filter: Since the inappropriate choice of the observation set would lead to privacy leakage of the private state, we design a randomized, optimization‐based filter that perturbs and then refines observations to maximize inference uncertainty of attackers.
+Finally, the proposed framework is demonstrated on a production–inventory case study, showing that our privacy filter significantly reduces the adversary’s capability to estimate the private production rate.
 1.4 Outline
 The rest of the paper is organized as follows.
-Section 2 introduces the system model and the inference attack.
-Section 3 defines the volumetric privacy measure, provides computational approaches for privacy level evaluation, and discusses the properties of the proposed measure.
+Section 2 introduces the system model, the inference attack and defines the volumetric privacy and utility measure.
+Section 3 provides computational approaches for privacy level evaluation, and discusses the properties of the proposed measure.
 Section 4 presents an optimal privacy filter design to mitigate privacy leakage while maintaining a certain utility level.
@@
 We use italic letters to denote the set of unknown variables, e.g., $\mathcal{X}$ and $\mathcal{Y}$ for $X$ and $Y$.
-We use the vector $\left[\begin{array}[]{c}\underline{X}\\ \overline{X}\\ \end{array}\right]$ to describe the interval $\mathcal{X}$, i.e., $\mathcal{X}=\left\{X|\underline{X}\leqslant X\leqslant\overline{X}\right\}$, where $\underline{X}$ and $\overline{X}$ are the lower and upper endpoints, respectively.
-The interval $\mathcal{X}$ can also be represented with
-$\mathcal{X}=\left\{c^{x}+\mathrm{diag}\left(p^{x}\right)\alpha:\alpha\in\mathcal{R}^{n_{x}},\left|\alpha\right|_{\infty}\leqslant 1\right\},$
-where $c^{x}=\frac{\overline{X}+\underline{X}}{2}$ and $p^{x}=\frac{\overline{X}-\underline{X}}{2}$ are the center point and radius of the interval.
-Besides, given the block matrix $A=\left[A_{1},A_{2}\right]$, the multiplication between the matrix $A$ and the interval $\mathcal{X}$ is defined as $A\mathcal{X}=A_{1}\underline{X}+A_{2}\overline{X}$.
-Also, if $\mathcal{X}$ and $\mathcal{Y}$ are intervals, then we use $\mathcal{X}+\mathcal{Y}$ to denote the interval $\left[\begin{array}[]{c}\underline{X}+\underline{Y}\\ \overline{X}+\overline{Y}\\ \end{array}\right]$, and use $\mathcal{X}-\mathcal{Y}$ to represent the interval $\left[\begin{array}[]{c}\underline{X}-\overline{Y}\\ \overline{X}-\underline{Y}\\ \end{array}\right]$.
-As for the non-interval set $\mathcal{Z}$, we use $A\mathcal{Z}$ to denote the new set $\left\{AZ|Z\in\mathcal{Z}\right\}$, and use $\mathcal{Z}\oplus\mathcal{R}$ to represent $\left\{Z+R|X\in\mathcal{Z},R\in\mathcal{R}\right\}$.
-Furthermore, the 1-norm of the column vector $b$ with n dimensions is defined as $\left\|b\right\|_{1}=\sum_{i=1}^{n}\left|b\left(i\right)\right|$, and $b^{\top}$ is the transpose of $b$.
+For the non-interval set $\mathcal{Z}$, we use $A\mathcal{Z}$ to denote the set $\left\{AZ|Z\in\mathcal{Z}\right\}$, and use $\mathcal{Z}\oplus\mathcal{R}$ to represent $\left\{Z\!+\!R|Z\in\mathcal{Z},R\in\mathcal{R}\right\}$.
+Furthermore, the 1-norm of the column vector $b$ with n dimensions is defined as $\left\|b\right\|_{1}=\sum_{i=1}^{n}\left|b\left(i\right)\right|$ with the absolute value $\left|b\left(i\right)\right|$, and $b^{\top}$ is the transpose of $b$.
+The 1-norm of the matrix $A$ is defined as $\left\|A\right\|_{1}\!\frac{\Delta}{=}\!\sum_{i,j}^{n,m}{\left|a_{i,j}\right|}$.
 The vector $\mathbf{1}_{n_{x}}$ denotes a column vector of ones with $n_{x}$ dimensions, while $I_{n_{x}\!\times\!n_{x}}$ represents an identity matrix of size $n_{x}\times n_{x}$.
+The operator $diag(v)$ denotes a diagonal matrix constructed from the vector $v$.
 2 System Model and Inference Attack
@@
 $\mathbf{G_{1}}:\left\{\begin{array}[]{c}X_{k}=A_{1}X_{k-1}+A_{2}Y_{k-1}+B_{1}W_{k}^{x}\\ Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+B_{2}W_{k}^{y}\\ \end{array}\right$., (3)
-where $A_{1}$ and $A_{2}$ are invertible, $Y_{k}\in\mathcal{R}^{n}$ is the private state, $X_{k}\in\mathcal{R}^{n}$ is the public state to be released, $W_{k}^{x}\in\mathcal{W}_{k}^{x}\subseteq\mathcal{R}^{m}$ and $W_{k}^{y}\in\mathcal{W}_{k}^{y}\subseteq\mathcal{R}^{m}$ are the UBB noises.
-Also, the initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively.
-When $A_{3}$ is zero, the model $\mathbf{G_{1}}$ can be regarded as a dynamic system with the public state $X_{k}$ and the private input $Y_{k}$.
-For analysis convenience, we assume that $\mathcal{X}_{0|-1}$, $\mathcal{Y}_{0|-1}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$ are intervals.
-For other types of bounded sets, such as zonotopes [14] and ellipses [4], the analysis is left for future investigation.
-Note that we do not assume the prior knowledge of probability distributions of the process noises, e.g., Gaussian distribution, which is different from most existing literature in privacy protection for dynamic systems.
-Furthermore, we assume that the adversary has full knowledge of system model $\mathbf{G_{1}}$ and will collect information of the public state to infer the private state.
-2.2 Motivating Example
-Consider the problem of privacy leakage related to the production rate in inventory control systems.
-To enhance sales, companies often need to share inventory levels with distributors.
-However, inventory levels are closely correlated with the production rate, which serves as a key indicator of the factory’s production strategy, operational efficiency, and supply chain dynamics.
-Therefore, the production rate constitutes highly sensitive information.
-If accessed by an adversary, this data could be exploited to develop competitive strategies, potentially compromising the company’s market position and operational integrity.
-The inventory of production can be modeled as follows [17, 29],
-$X_{k}=X_{k-1}+Y_{k-1}-W_{k}^{x},$ (4)
-where $X_{k}$ is the inventory, $Y_{k}$ is the production rate, and $W_{k}^{x}$ is the uncertain demand.
-The production rate is usually chosen to maintain the inventory around a desired level, which can be modeled as follows,
-$Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+W_{k}^{y},$ (5)
-where the uncertain process noise $W_{k}^{y}$ is due to the adversary’s limited knowledge about the company’s production capability.
-When the uncertainty of demand and production rate is small, e.g., $W_{k}^{x}=D^{x}$ and $W_{k}^{y}=D^{y}$ are constant, then the adversary can approximately reconstruct the production rate with the inventory level via
-$Y_{k-1}\approx X_{k}-X_{k-1}+D^{x},$ (6)
-and
-$Y_{k}\approx A_{3}X_{k-1}+A_{4}Y_{k-1}+D^{y},$ (7)
-which causes privacy leakage and increases potential risks of business competition.
-However, in practical scenarios, the uncertainty associated with demand and production capability is bounded but cannot be ignored.
-Inferring the private production rate, therefore, becomes a complex but still feasible task.
-Due to the presence of unknown noise terms that belong to bounded sets without probabilistic assumptions, multiple private states may correspond to the same public state with equal probability, leading to the formation of an uncertainty set.
-To address this challenge, we define the adversary’s inference attack on the private state using set operations in the following sections.
+where $A_{1}$ and $A_{2}$ are invertible, $Y_{k}\in\mathcal{R}^{n}$ is the private state, $X_{k}\in\mathcal{R}^{n}$ is the public state to be released, $W_{k}^{x}\in\mathcal{W}_{k}^{x}\subseteq\mathcal{R}^{m}$ and $W_{k}^{y}\in\mathcal{W}_{k}^{y}\subseteq\mathcal{R}^{m}$ are the unknown disturbance with bounded sets $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$.
+Note that the underlying probability distributions of $W_{k}^{x}$ and $W_{k}^{y}$ are unknown.
+Besides, the initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively.
+We assume that the adversary has full knowledge of system model $\mathbf{G_{1}}$ and will collect information of the public state to infer the private state.
+2.2 Motivating Examples
+In this subsection, we consider two motivating examples to illustrate the necessity of protecting privacy of systems $\mathbf{G_{1}}$.
+Production-inventory system: In supply chain management [28, 29], the inventory level $X_{k}$ and the production rate $Y_{k}$ evolves according to $\mathbf{G}_{1}$.
+While firms may disclose inventory information $X_{k}$ to distributors to boost sales, the production rate $Y_{k}$ contains sensitive strategic information such as production efficiency and supply chain operations.
+Since $X_{k}$ and $Y_{k}$ are correlated, releasing $X_{k}$ directly risks revealing private production details.
+Therefore, it is necessary to transform or mask observations to preserve the privacy of $Y_{k}$ while maintaining the utility of public inventory data $X_{k}$.
+Traffic management system: In intelligent transportation, vehicles may report their velocities to a central controller to optimize traffic flow, e.g., by adjusting the speed limit on highways.
+Given bounded disturbances from environmental factors like uneven ground, the vehicle dynamics fit the model $\mathbf{G}_{1}$ with unknown-but-bounded disturbance.
+Here, velocity $X_{k}$ can be considered public data used for traffic management, while position $Y_{k}$ is private, as it can be used to identify individual vehicles.
+To protect location privacy, vehicles may intentionally report blurred or randomized velocity observations that preserve system utility but reduce the risk of precise location inference.
 2.3 Inference Attack
-We assume that the adversary observes a set of public states, denoted as $\mathcal{M}_{k|k}^{x}$, which includes the actual public state value $X_{k}=x_{k}$.
-Also, every element of $\mathcal{M}_{k|k}^{x}$ is considered a potential candidate for the actual public state.
-Notably, if the public state is directly transmitted to the adversary without any processing, the public state set $\mathcal{M}_{k|k}^{x}$ contains only the actual public state.
+Due to the presence of unknown disturbance terms belonging to bounded sets, multiple private states may correspond to the same public state.
+Consequently, for a given public state set, an adversary cannot determine the exact private state; instead, it can only identify a corresponding uncertainty set.
+we next define how the adversary infers the private state using set-theoretic operations.
+We assume that the adversary observes a set of public states, denoted as $\mathcal{M}_{k|k}^{x}$, which includes the actual public state value $X_{k}=x_{k}$ and other elements to obfuscate the adversary’s estimation of private states.
 Based on the observed public state set, the adversary identifies all potential values of the private state that align with $\mathcal{M}_{k|k}^{x}$ to construct its uncertainty set.
@@
 We next define the inference attack recursively.
-At time $k$, given the public state set $\mathcal{X}_{k-1|k-1}$ and the uncertainty private state set $\mathcal{Y}_{k-1|k-1}$, the set of public state can be predicted based on the system model (3), i.e.,
-$\mathcal{X}_{k|k-1}=A_{1}\mathcal{X}_{k-1|k-1}\oplus A_{2}\mathcal{Y}_{k-1|k-1}\oplus B_{1}\mathcal{W}_{k}^{x}.$ (8)
-Similarly, the set of private state can be predicted via
+At time $k$, given the public state set $\mathcal{X}_{k-1|k-1}$ and the uncertainty private state set $\mathcal{Y}_{k-1|k-1}$, the set of states can be predicted based on the system model (3), i.e.,
+$\mathcal{X}_{k|k-1}$ $=A_{1}\mathcal{X}_{k-1|k-1}\oplus A_{2}\mathcal{Y}_{k-1|k-1}\oplus B_{1}\mathcal{W}_{k}^{x},$ (4)
 $\mathcal{Y}_{k|k-1}$ $=A_{3}\mathcal{X}_{k-1|k-1}\oplus A_{4}\mathcal{Y}_{k-1|k-1}\oplus B_{2}\mathcal{W}_{k}^{y}.$ (5)
 After receiving the observation set of the public state $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$, the adversary extracts new information from $\mathcal{M}^{x}_{k|k}$ and updates the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via the following steps,
-$\mathcal{M}_{k-1|k}^{x}=$$\left\{M^{x}_{k-1|k}|X_{k|k}=A_{1}M^{x}_{k-1|k}+A_{2}Y_{k-1|k-1}+B_{1}W_{k}^{x},\right.$
-$\forall X_{k|k}$ $\in\mathcal{M}^{x}_{k|k},\left.\forall Y_{k-1|k-1}\in\mathcal{Y}_{k-1|k-1},\forall W_{k}^{x}\in\mathcal{W}_{k}^{x}\right\}\!,$ (10)
-$\mathcal{M}_{k-1|k}^{y}=$$\left\{M^{y}_{k-1|k}|X_{k|k}=A_{1}X_{k-1|k}+A_{2}M^{y}_{k-1|k}+B_{1}W_{k}^{x},\right.$
-$\forall X_{k|k}$ $\in\mathcal{M}^{x}_{k|k},\left.\forall X_{k-1|k-1}\in\mathcal{X}_{k-1|k-1},\forall W_{k}^{x}\in\mathcal{W}_{k}^{x}\right\}\!,$ (11)
+$\!\!\!\!\mathcal{M}_{k\!-\!1|k}^{x}\!\!=\!A_{1}^{\!-\!1}\!\mathcal{M}_{k|k}^{x}$ $\!\oplus\!\left(\!-\!A_{1}^{\!-\!1}\!A_{2}\!\right)\!\mathcal{Y}_{k\!-\!1|k\!-\!1}\!\oplus\!\left(-\!A_{1}^{\!-\!1}\!B_{1}\!\right)\!\mathcal{W}_{k}^{x}\!,\!\!$ (6)
+$\!\!\!\!\mathcal{M}_{k\!-\!1|k}^{y}\!=\!A_{2}^{\!-\!1}\!\mathcal{M}_{k|k}^{x}$ $\!\oplus\!\left(\!-A_{2}^{\!-\!1}\!A_{1}\!\right)\!\mathcal{X}_{k\!-\!1|k\!-\!1}\!\oplus\!\left(\!-A_{2}^{\!-\!1}\!B_{1}\!\right)\!\mathcal{W}_{k}^{x}\!,\!\!$ (7)
 $\mathcal{X}_{k-1|k}$ $=\mathcal{M}_{k-1|k}^{x}\cap\mathcal{X}_{k-1|k-1},$ (8)
 $\mathcal{Y}_{k-1|k}$ $=\mathcal{M}_{k-1|k}^{y}\cap\mathcal{Y}_{k-1|k-1},$ (9)
-where it first computes the possible sets of the public and private states, i.e., $\mathcal{M}_{k-1|k}^{x}$ and $\mathcal{M}_{k-1|k}^{y}$, based on the system model (3) and the observation $\mathcal{M}^{x}_{k|k}$ in (2.3) and (2.3), and then reduces the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via intersection operations in (12) and (13).
+where it first computes the possible sets of the public and private states, i.e., $\mathcal{M}_{k-1|k}^{x}$ and $\mathcal{M}_{k-1|k}^{y}$, based on the system model (3) and the observation $\mathcal{M}^{x}_{k|k}$ in (6) and (7), and then reduces the uncertainty sets of $X_{k-1}$ and $Y_{k-1}$ via intersection operations in (8) and (9).
 According to the system dynamics (3), the adversary estimates the uncertainty set of $Y_{k}$ via the following forward inference,
 $\mathcal{Y}_{k|k}=A_{3}\mathcal{X}_{k-1|k}\oplus A_{4}\mathcal{Y}_{k-1|k}\oplus B_{2}\mathcal{W}_{k}^{y}.$ (10)
-Finally, the public state set can be further calibrated via intersection,
+Finally, the public state set can be further calibrated via,
 $\mathcal{X}_{k|k}=\mathcal{M}^{x}_{k|k}\cap\mathcal{M}^{x}_{k|k-1},$ (11)
-where
 $\mathcal{M}^{x}_{k|k-1}=$ $A_{1}\mathcal{X}_{k-1|k}\oplus A_{2}\mathcal{Y}_{k-1|k}\oplus B_{1}\mathcal{W}_{k}^{x},$ (12)
-is the predicted uncertainty set of $X_{k}$ based on the calibrated sets $\mathcal{X}_{k-1|k}$ and $\mathcal{Y}_{k-1|k}$.
-Starting from $k=0$, with the initial uncertainty sets $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, the adversary can recursively update the uncertainty sets of $X_{k}$ and $Y_{k}$ via the backward calibration (2.3)-(13), and the forward inference (14)-(16).
-The backward calibration (2.3)-(13) reduces the uncertainty of $X_{k-1}$ and $Y_{k-1}$, which leads to the following proposition.
+where $\mathcal{M}^{x}_{k|k-1}$ is the predicted uncertainty set of $X_{k}$ based on the calibrated sets $\mathcal{X}_{k-1|k}$ and $\mathcal{Y}_{k-1|k}$.
+Starting from $k=0$, with the initial uncertainty sets $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, the adversary can recursively update the uncertainty sets of $X_{k}$ and $Y_{k}$ via the backward calibration (6)-(9), and the forward inference (10)-(12).
+The backward calibration (6)-(9) reduces the uncertainty of $X_{k-1}$ and $Y_{k-1}$, which leads to the following proposition.
 Proposition 1.
-For any $k\geqslant 1$, the adversary’s uncertainty private state set (14) is a subset of its prediction set (9).
-Also, given the uncertainty sets $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$ containing the actual system states $X_{k-1}=x_{k-1}$ and $Y_{k-1}=y_{k-1}$, if the observation set $\mathcal{M}^{x}_{k|k}$ contains the actual public state $X_{k}=x_{k}$, then the inference result $\mathcal{Y}_{k|k}$ also contains the actual private state $Y_{k}=y_{k}$.
+For any $k\geqslant 1$, the adversary’s uncertainty set for the private state (10) is a subset of its corresponding prediction set (5), i.e., $\mathcal{Y}_{k|k}\subseteq\mathcal{Y}_{k|k-1}$.
+Moreover, given uncertainty sets $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$ that contain the true system states $X_{k-1}\!=\!x_{k-1}$ and $Y_{k-1}\!=\!y_{k-1}$, if the observation set $\mathcal{M}^{x}_{k|k}$ contains the true public state $X_{k}\!=\!x_{k}$, then the inference set $\mathcal{Y}_{k|k}$ contains the true private state $Y_{k}\!=\!y_{k}$.
 Proof 2.1.
-The proof is omitted due to the simplicity.
-According to Proposition 1, the inference attack results in a smaller uncertainty private state set $\mathcal{Y}_{k|k}$ that contains the actual private state $y_{k}$, meaning that there are fewer possible private states corresponds to the same observation set $\mathcal{M}_{k}^{x}$.
-In particular, if the uncertainty set $\mathcal{Y}_{k|k}$ contains only one element, then the adversary can obtain actual private state.
-In conclusion, the adversary can reduce its uncertainty of the private state via the inference attack.
-2.4 Privacy Measure and Defense
-The aforementioned inference attack relies on set operations, and its outcome is the uncertainty set of the private state, denoted as $\mathcal{Y}_{k|k}$.
-The first key question addressed in this paper is how to evaluate the privacy level of the private state using the uncertainty set $\mathcal{Y}_{k|k}$.
-Also, set operations, such as the addition in (2.3) and the intersection in (12), involve all possible combinations of elements from different collections.
-These operations are computationally challenging due to the continuous nature of the state space.
-Consequently, it is essential to provide efficient computational tools for privacy level evaluation.
-Moreover, from a defense perspective, another critical task is the design of effective mechanisms to mitigate privacy leakage.
-To address these challenges, we propose a volumetric privacy measure and its computation approach in Section 3.
-Furthermore, we present an optimal privacy filter design to reduce privacy leakage in Section 4.
-3 Volumetric Privacy Measure
-In this section, we define the volume of uncertainty sets as a quantitative measure of the privacy level.
-We then introduce an interval-based approach to approximate the computation of state uncertainty sets and evaluate the corresponding privacy levels.
-Additionally, we analyze the properties of the volumetric measure, showing that both the privacy level and the uncertainty reduction are bounded by the new information derived from the observation set.
+Since $\mathcal{M}^{x}_{k|k}$ contains $x_{k}$, it follows from (6) that $x_{k-1}\!\in\!\mathcal{M}^{x}_{k-1|k}$.
+As $x_{k-1}$ also belongs to $\mathcal{X}_{k-1|k-1}$, their intersection $\mathcal{X}_{k-1|k}$ necessarily contains $x_{k-1}$.
+By the same reasoning, $y_{k-1}\!\in\!\mathcal{Y}_{k-1|k}$.
+Propagating through the system dynamics $\mathbf{G}_{1}$ yields $y_{k}\!\in\!\mathcal{Y}_{k|k}$.
+Finally, since $\mathcal{X}_{k-1|k}\!\subseteq\!\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k}\!\subseteq\!\mathcal{Y}_{k-1|k-1}$, we have $\mathcal{Y}_{k|k}\!\subseteq\!\mathcal{Y}_{k|k-1}$.
+According to Proposition 1, the adversary can reduce its uncertainty of the private state via the inference attack since it can obtain a smaller uncertainty private state set $\mathcal{Y}_{k|k}$ that contains the actual private state $y_{k}$.
+In particular, if the uncertainty set $\mathcal{Y}_{k|k}$ contains only one element, then the adversary can obtain the actual private state.
 2.4 Privacy and Utility Measures
 As discussed in Section 2.3, the uncertainty set of private state $\mathcal{Y}_{k|k}$ encompasses all possible elements that correspond to the same observation set, $\mathcal{M}_{k|k}^{x}$.
-However, due to the continuous nature of the state space, the number of elements in such sets is uncountable.
-To address this, we propose using the volume of the uncertainty set as a quantitative measure of privacy, defined as follows,
-$P_{k}\left(\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right),$ (17)
-where $\mathrm{Vol}\left(\cdot\right)\geqslant 0$ is a Lebesgue measure.
-When $\mathcal{Y}_{k|k}$ is an interval, we can compute its volume via $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)=\sum_{i=1}^{n}{\left(\overline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k}\left(i\right)\right)}$, where $\overline{Y}_{k|k}$ and $\underline{Y}_{k|k}$ are the upper and lower endpoints of the interval $\mathcal{Y}_{k|k}$, respectively.
-Since the uncertainty set $\mathcal{Y}_{k|k}$ contains the actual value of private state $Y_{k}=y_{k}$, specially, the adversary can accurately access to $y_{k}$ if the volume of $\mathcal{Y}_{k|k}$ is zero.
-On the other hand, if the volume $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is large, then the range of private state will be large, which increases the difficulties of selecting the correct value as the estimation result.
-Therefore, the volume of $\mathcal{Y}_{k|k}$ describes the amount of the adversary’s inference uncertainty about the private state as well as the privacy level.
-We can increase the value of $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ to increase the privacy level to protect the system from inference attack.
-Besides, since the uncertainty set $\mathcal{Y}_{k|k}$ is a subset of its predicted version $\mathcal{Y}_{k|k-1}$, we have
-$\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\leqslant\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right),$ (18)
-meaning that the adversary’s inference uncertainty about the private state is bounded by its prior knowledge $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$.
-In other words, the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ cannot exceed $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$.
-Furthermore, we propose to use the following measure to measure the utility of public state set $\mathcal{X}_{k|k}$,
-$U_{k}\left(\mathcal{X}_{k|k}\right)=\frac{1}{\mathrm{Vol}\left(\mathcal{X}_{k|k}\right)},$ (19)
-since a larger $\mathcal{X}_{k|k}$ provides more possible values of the public state to the receiver, i.e., increases the estimation error.
-Consequently, a larger value of $U_{k}\left(\mathcal{X}_{k|k}\right)$ indicates a higher data utility of the public state set.
-3.2 Inference Attack Approximation
-In this subsection, we investigate the inference approximation for multi-dimensional systems using the tightest interval method [1].
-This approach simplifies the computation of the privacy measure and provides the foundation for designing the optimal privacy filter as discussed in Section 4.
+The attacker has more inference uncertainty about the private state if $\mathcal{Y}_{k|k}$ contains more elements.
+However, since the state space is continuous, the number of elements in such sets is uncountable.
+To address this, we propose to use the volume of the uncertainty set as a quantitative measure of privacy.
+Specifically, we define
+$\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)=\int_{\mathcal{Y}_{k|k}}\mathrm{d}y,$ (13)
+where $\mathrm{Vol}(\mathcal{Y}_{k|k})$ denotes the Lebesgue measure (i.e., the geometric volume) of the set $\mathcal{Y}_{k|k}\subseteq\mathcal{R}^{n}$.
+We then define the privacy measure at time $k$ as
+$P_{k}\left(\mathcal{Y}_{k|k}\right):=\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right).$ (14)
+Since the uncertainty set $\mathcal{Y}_{k|k}$ contains the actual value of private state $Y_{k}=y_{k}$, the adversary can accurately access to $y_{k}$ if the volume of $\mathcal{Y}_{k|k}$ is zero.
+To protect the system from inference attack, we can increase the privacy level of the private state via maximizing $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$.
+On the other hand, it causes more utility distortion of the public state if the observation set $\mathcal{M}^{x}_{k|k}$ contains more elements, since it is more difficult for the receiver to recover the actual public state.
+To address this, we would reduce distortion via maximizing the following defined utility of public state set $\mathcal{M}^{x}_{k|k}$,
+$U_{k}\left(\mathcal{M}^{x}_{k|k}\right)=1/\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right).$ (15)
+Notably, set operations in the proposed inference attack and the volume computation are generally computational costly due to the continuous nature of the state space.
+Consequently, it is essential to provide efficient computational tools for privacy level evaluation.
+Also, from a defense perspective, another critical task is the design of effective mechanisms to mitigate privacy leakage.
+To address these challenges, we study computation approaches for inference attack in Sec.
+3, and present an optimal privacy filter design to reduce privacy leakage in Sec.
+4.
+3 Inference Attack Approximation
+As addressed in existing set-membership estimation approaches [18, 19], the set operations involved in inference attacks can be computationally expensive.
+To mitigate this complexity, uncertainty sets are often restricted to specific geometric forms, enabling more efficient implementation of set operations.
+However, more complex representations generally incur higher computational costs in volume evaluation.
+In the following, we present two approximation methods for implementing inference attacks and analyze their computational complexity, based on which we establish properties of volumetric privacy.
+3.1 Inference Attack Approximation via CCGs
+In this subsection, we show that the inference attack can be approximated using the constrained convex generator (CCG), a general set representation proposed in [30].
+Definition 2.
+[30] The constrained convex generator $\mathcal{Z}=\left(G,c,A,b,\mathcal{C}\right)\subset\mathcal{R}^{n}$ is defined as
+$\mathcal{Z}=\left\{G\xi+c:A\xi=b,\xi\in\mathcal{C}\right\},$ (16)
+where $G\in\mathcal{R}^{n\times n_{g}}$, $c\in\mathcal{R}^{n}$, $A\in\mathcal{R}^{n_{c}\times n_{g}}$, $b\in\mathcal{R}^{n_{c}}$ and $\mathcal{C}=\left\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{n_{p}}\right\}$, and $\mathcal{C}_{i}\subset\mathcal{R}^{m_{i}}$ are convex sets with $\sum_{i=1}^{n_{p}}{m_{i}}=n_{g}$.
+The CCG encompasses a wide range of useful set representations, including zonotopes, ellipsoids, and intervals [30].
+Moreover, common set operations such as the Minkowski sum and intersection admit analytical expressions, enabling its application to approximate the inference attack.
+Proposition 3.
+[30] Given CCGs $\mathcal{X}\!=\!\left(G_{x},c_{x},A_{x},b_{x},\mathcal{C}_{x}\right)\!\subset\!\mathcal{R}^{n}$ and $\mathcal{Y}\!=\!\left(G_{y},c_{y},A_{y},b_{y},\mathcal{C}_{y}\right)\!\subset\!\mathcal{R}^{n}$, and a matrix $R\in\mathcal{R}^{m\times n}$, we have
+$R\mathcal{X}=\left(RG_{x},Rc_{x},A_{x},b_{x},\mathcal{C}_{x}\right),$
+$\mathcal{X}\!\oplus\!\mathcal{Y}$ $\!=\!\left(\!\left[\begin{matrix}G_{x}&\!G_{y}\\ \end{matrix}\right]\!,c_{x}\!+\!c_{y},\mathrm{diag}\left(\!\left[\begin{matrix}A_{x}&\!A_{y}\\ \end{matrix}\right]\!\right)\!,\left[\!\begin{array}[]{c}b_{x}\\ b_{y}\\ \end{array}\!\right]\!,\left\{\mathcal{C}_{x},\mathcal{C}_{y}\right\}\!\!\right)\!,$ (19)
+$\mathcal{X}\cap\mathcal{Y}$ $=\left(\left[\begin{matrix}G_{x}&0\\ \end{matrix}\right],c_{x},\left[\begin{matrix}A_{x}&0\\ 0&A_{y}\\ G_{x}&-G_{y}\\ \end{matrix}\right],\left[\begin{array}[]{c}b_{x}\\ b_{y}\\ c_{y}-c_{x}\\ \end{array}\right],\left\{\mathcal{C}_{x},\mathcal{C}_{y}\right\}\right).$ (23)
+According to the computation rules in Proposition 3, the inference attack from (6) to (12) can be directly implemented, if we assume that the uncertainty sets $\mathcal{W}^{x}_{k}$, $\mathcal{W}^{y}_{k}$, $\mathcal{X}_{0|-1}$, and $\mathcal{Y}_{0|-1}$ are represented as CCGs.
+However, as shown in the next lemma, the computational complexity of CCG-based inference grows exponentially over time.
+Proposition 4.
+The computational complexity of the CCG-based inference attack at time $k$ is at least $\mathcal{O}\!\left(c^{k-1}n^{3}\right)$ for some constant $c>1$, and both the column dimension of the generator matrices $G$ and the number of constraints grow exponentially with $k$.
+Proof 3.1.
+The dominant operation in the inference attack from (6) to (12) is the multiplication of an $n\times n$ matrix with an $n\times m$ matrix, which has computational complexity $\mathcal{O}(mn^{2})$.
+For simplicity, we assume that at time $k$ the sets $\mathcal{X}_{k-1|k-1}$, $\mathcal{Y}_{k-1|k-1}$, $\mathcal{W}^{x}_{k}$, $\mathcal{W}^{y}_{k}$, and $\mathcal{M}^{x}_{k|k}$ all have generator matrices $G$ of size $n\times n$ and are described by $n$ constraints.
+According to Proposition 3, after one inference step, the Minkowski sum and intersection operations cause the generator matrix in $\mathcal{X}_{k|k}$ to grow to size $n\times(c\,n)$ for some constant $c>1$, while the number of constraints increases by a factor $d>1$.
+Thus, both the column dimension of $G$ and the number of constraints grow exponentially with $k$.
+Consequently, due to the exponentially increasing column dimension, the computational complexity of matrix multiplications in the inference attack at time $k$ is at least $\mathcal{O}(c^{k-1}n^{3})$.
+The high computational complexity of CCG-based inference renders real-time implementation of the inference attack and the privacy filter design in Sec.
+4 intractable for large $k$.
+Order-reduction techniques can be employed to reduce this complexity, albeit at the cost of some loss in inference accuracy.
+However, such techniques also complicate the analysis of the proposed volumetric privacy metric.
+For clarity and focus, we defer a detailed discussion of these techniques to future work.
+3.2 Inference Attack Approximation via Interval Analysis
+We next consider an interval-based approximation approach for computing the inference sets.
+This approach is a special case of CCG-based inference but significantly reduces both computational and analytical complexity due to the efficiency of interval arithmetic.
+Moreover, interval-based inference serves as the foundation for designing the optimal privacy filter, as discussed in Sec.
+4.
+Definition 5.
+The interval $\mathcal{X}=\{X\mid\underline{X}\leqslant X\leqslant\overline{X}\}$ is defined as $\begin{bmatrix}\underline{X}\\[3.00003pt] \overline{X}\end{bmatrix}$ with the lower and upper bounds $\underline{X}$ and $\overline{X}$.
+An interval $\mathcal{X}$ can also be expressed as a special case of the CCG representation $\mathcal{X}=\left\{\mathrm{diag}(p^{x})\,\xi+c^{x}\;:\;\xi\in\mathcal{R}^{n_{x}},\;\|\xi\|_{\infty}\leqslant 1\right\}$, with the center point $c^{x}=\frac{\overline{X}+\underline{X}}{2}$ and the radius $p^{x}=\frac{\overline{X}-\underline{X}}{2}$.
+Also, the volume of the interval $\mathcal{X}$ is computed as $\mathrm{Vol}\left(\mathcal{X}\right)=\sum_{i=1}^{n}{\left(\overline{X}\left(i\right)-\underline{X}\left(i\right)\right)}$ where $\overline{X}(i)$ and $\underline{X}(i)$ denote the upper and lower bounds of the $i$-th dimension of the interval $\mathcal{X}$, respectively.
+Given a block matrix $A=[A_{1},\,A_{2}]$, the multiplication of $A$ with an interval $\mathcal{X}$ is defined as $A\mathcal{X}=A_{1}\underline{X}+A_{2}\overline{X}$.
+If $\mathcal{X}$ and $\mathcal{Y}$ are intervals, their Minkowski sum is $\mathcal{X}\oplus\mathcal{Y}=\begin{bmatrix}\underline{X}+\underline{Y}\\[3.00003pt] \overline{X}+\overline{Y}\end{bmatrix}$, and their difference, used only for volume evaluation, is $\mathcal{X}\setminus\mathcal{Y}=\begin{bmatrix}\underline{X}-\underline{Y}\\[3.00003pt] \overline{X}-\overline{Y}\end{bmatrix}$.
+The intersection of $\mathcal{X}$ and $\mathcal{Y}$ is $\mathcal{X}\cap\mathcal{Y}=\begin{bmatrix}\max\{\underline{X},\underline{Y}\}\\[3.00003pt] \min\{\overline{X},\overline{Y}\}\end{bmatrix}$.
+We now assume that the uncertainty sets $\mathcal{W}^{x}_{k}$, $\mathcal{W}^{y}_{k}$, $\mathcal{X}_{0|-1}$, and $\mathcal{Y}_{0|-1}$ are represented as intervals.
+Under this assumption, the interval-based inference attack can be implemented using the following lemma.
 Lemma 6.
-The tightest recursive inference interval from (2.3) to (13) can be computed via
-$\mathcal{M}_{k-1|k}^{x}=$ $\Psi\left(A_{1}^{-1}\right)\mathcal{M}^{x}_{k|k}+\Psi\left(-A_{1}^{-1}A_{2}\right)\mathcal{Y}_{k-1|k-1}$
-$+\Psi\left(-A_{1}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (20)
-$\mathcal{M}_{k-1|k}^{y}=$ $\Psi\left(A_{2}^{-1}\right)\mathcal{M}^{x}_{k|k}+\Psi\left(-A_{2}^{-1}A_{1}\right)\mathcal{Y}_{k-1|k-1}$
-$+\Psi\left(-A_{2}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (21)
+The recursive inference interval from (6) to (9) can be computed via
+$\mathcal{M}_{k-1|k}^{x}$ $=\Psi\left(A_{1}^{-1}\right)\mathcal{M}^{x}_{k|k}\oplus\Psi\left(-A_{1}^{-1}A_{2}\right)\mathcal{Y}_{k-1|k-1}$
+$\oplus\Psi\left(-A_{1}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (24)
+$\mathcal{M}_{k-1|k}^{y}$ $=\Psi\left(A_{2}^{-1}\right)\mathcal{M}^{x}_{k|k}\oplus\Psi\left(-A_{2}^{-1}A_{1}\right)\mathcal{X}_{k-1|k-1}$
+$\oplus\Psi\left(-A_{2}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (25)
 $\mathcal{X}_{k-1|k}$ $=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k-1|k}^{x},\underline{X}_{k-1|k-1}\right\}\\ \min\left\{\overline{M}_{k-1|k}^{x},\overline{X}_{k-1|k-1}\right\}\\ \end{array}\right],$ (28)
 $\mathcal{Y}_{k-1|k}$ $=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k-1|k}^{y},\underline{Y}_{k-1|k-1}\right\}\\ \min\left\{\overline{M}_{k-1|k}^{y},\overline{Y}_{k-1|k-1}\right\}\\ \end{array}\right],$ (31)
-$\!\!\!\!\!\!\mathcal{M}_{k|k-1}^{x}\!=\!\Psi\!\left(A_{1}\right)\!\mathcal{X}_{k-1|k}\!+\!\Psi\!\left(A_{2}\right)\!\mathcal{Y}_{k-1|k}\!+\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x},$ (28)
+$\!\!\!\!\!\!\mathcal{M}_{k|k-1}^{x}\!=\!\Psi\!$ $\left(A_{1}\right)\!\mathcal{X}_{k-1|k}\!\oplus\!\Psi\!\left(A_{2}\right)\!\mathcal{Y}_{k-1|k}\!\oplus\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x},$ (32)
 $\mathcal{X}_{k|k}$ $=\left[\begin{array}[]{c}\max\left\{\underline{M}_{k|k}^{x},\underline{M}_{k|k-1}^{x}\right\}\\ \min\left\{\overline{M}_{k|k}^{x},\overline{M}_{k|k-1}^{x}\right\}\\ \end{array}\right],$ (35)
-$\!\!\!\!\!\!\!\!\mathcal{Y}_{k|k}=\Psi\left(A_{3}\right)\mathcal{X}_{k-1|k}+\Psi\left(A_{4}\right)\mathcal{Y}_{k-1|k}+\Psi\left(B_{2}\right)\mathcal{W}_{k}^{y},$ (32)
+$\!\!\!\!\!\!\!\!\mathcal{Y}_{k|k}\!=\!\Psi\!$ $\left(A_{3}\right)\!\mathcal{X}_{k-1|k}\!\oplus\!\Psi\!\left(A_{4}\right)\mathcal{Y}_{k-1|k}\!\oplus\!\Psi\!\left(B_{2}\right)\mathcal{W}_{k}^{y}\!,\!$ (36)
 with
 $\Psi\left(\star\right)=\left[\begin{matrix}\frac{\star+\left|\star\right|}{2}&\frac{\star-\left|\star\right|}{2}\\ \frac{\star-\left|\star\right|}{2}&\frac{\star+\left|\star\right|}{2}\\ \end{matrix}\right].$
-Also, the tightest prior inference set of $Y_{k}$ is
-$\!\!\!\!\!\!\!\mathcal{Y}_{k|k-1}\!=\!\Psi\left(\!A_{3}\!\right)\!\mathcal{X}_{k-1|k-1}\!+\!\Psi\left(\!A_{4}\!\right)\!\mathcal{Y}_{k-1|k-1}\!+\!\Psi\left(\!B_{2}\!\right)\!\mathcal{W}_{k}^{y},$ (33)
+Also, the prior inference set of $Y_{k}$ is
+$\!\!\!\!\!\mathcal{Y}_{k|k\!-\!1}\!=\!\Psi\!\left(\!A_{3}\!\right)\!\mathcal{X}_{k-1|k-1}\!\oplus\!\Psi\!\left(\!A_{4}\!\right)\!\mathcal{Y}_{k-1|k-1}\!\oplus\!\Psi\!\left(\!B_{2}\!\right)\!\mathcal{W}_{k}^{y}\!,\!\!$ (37)
 if $k\geqslant 1$. If $k=0$, then $\mathcal{Y}_{0|0}=\mathcal{Y}_{0|-1}$ and
 $\mathcal{X}_{0|0}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{0|0}^{x},\underline{X}_{0|-1}\right\}\\ \min\left\{\overline{M}_{0|0}^{x},\overline{X}_{0|-1}\right\}\\ \end{array}\right].$ (40)
-Proof 3.1.
+Proof 3.2.
 See Appendix A.
-According to Lemma 2, the inference attack can be approximately computed using intervals.
-It can be verified that the intervals computed in Lemma 2 are tight when $X_{k}$ and $Y_{k}$ are scalar variables.
-3.3 Properties of the Inference Attack
+Given the interval-based inference approach described in Lemma 6, the computational complexity of the inference attack can be characterized as follows.
+Proposition 7.
+The computational complexity of the inference attack via interval analysis is $\mathcal{O}(n^{3})$.
+Proof 3.3.
+The dominant operation in the inference attack involves matrix multiplication.
+Since the matrix $\Psi(\star)$ has dimensions $(2n\times 2n)$, the corresponding computational complexity is $\mathcal{O}(n^{3})$.
+Although the matrix multiplication with large $n$ can still be computationally demanding, the complexity of interval-based inference is substantially lower and remains constant over time, in sharp contrast to the exponentially growing complexity of CCG-based inference.
+3.3 Properties of the Interval Inference Attack
 The inference attack exhibits certain properties.
-For instance, the radius of the uncertainty $\mathcal{Y}_{k|k}$, i.e., $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$ is bounded by a function of the radius of the process noise and the observation set, as stated below.
+For instance, the radius of the uncertainty $\mathcal{Y}_{k|k}$, i.e., $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$ is bounded by a function of the radius of the disturbance and the observation set, as stated below.
 Lemma 8.
@@
 $+\left|A_{4}\right|\left|A_{2}^{-1}B_{1}\right|p^{w,x}_{k}+\left|B_{2}\right|p^{w,y}_{k},$ (41)
-where $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$ and $p^{w,y}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$, respectively.
-Proof 3.2.
+where $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$ and $p^{w,y}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$, respectively, and $|A|$ is a matrix where each element is the absolute value of the corresponding element in $A$, i.e., $|A|=\left[|a_{i,j}|\right]$.
+Proof 3.4.
 See Appendix B.
-Since the volume of $\mathcal{Y}_{k|k}$ is the sum of $p_{k|k}^{y}$, the volume $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is also bounded by a function of $\overline{p}^{x}$.
-As a result, if the value of elements in $\overline{p}^{x}$ are small, i.e., if the observation set $\mathcal{M}_{k|k}^{x}$ is small, then the privacy level is low, and the adversary experiences less uncertainty after performing the inference attack.
+Since the volume of $\mathcal{Y}_{k|k}$ is the sum of $p_{k|k}^{y}$, $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is also bounded by a function of the radius of the observation set $\mathcal{M}_{k|k}^{x}$.
+Consequently, if $\mathcal{M}_{k|k}^{x}$ is small, then the privacy level would be low, and the adversary retains little uncertainty after performing the inference attack.
 Furthermore, by comparing the predicted and posterior uncertainty sets, as given by (5) and (10), the amount of uncertainty reduction can be quantified by $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, where
-$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\diagdown\mathcal{Y}_{k|k}.$
-Proposition 4.
-If $\mathcal{Y}_{k|k}$ is a subset of $\mathcal{Y}_{k|k-1}$, then the volume of $\Delta\mathcal{Y}_{k|k}$ is
-$\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\!.$
-Proof 3.3.
-Since $\mathcal{Y}_{k|k}$ is the subset of $\mathcal{Y}_{k|k-1}$, the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is equal to $\left\|\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}\right\|_{1}$, which can be computed as follows,
-$\left\|\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}\right\|_{1}$
-$=$ $\sum_{i=1}^{n_{y}}\left(\overline{Y}_{k|k-1}\left(i\right)-\overline{Y}_{k|k}\left(i\right)+\underline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)$
-$=$ $\sum_{i=1}^{n_{y}}{\left(\overline{Y}_{k|k-1}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)}-\sum_{i=1}^{n_{y}}{\left(\overline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k}\left(i\right)\right)}$
-$=$ $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right),$
-where $\left(\overline{Y}_{k|k-1}\left(i\right)-\overline{Y}_{k|k}\left(i\right)+\underline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)$ is the length of i-th interval of $\Delta\mathcal{Y}_{k|k}$.
-According to Proposition 4, to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, it is equivalent to reduce the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ since the amount of prior uncertainty $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$ is fixed at time $k$.
+$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\setminus\mathcal{Y}_{k|k},$
+is the difference between $\mathcal{Y}_{k|k-1}$ and $\mathcal{Y}_{k|k}$.
+Since $\mathcal{Y}_{k|k}$ is a subset of $\mathcal{Y}_{k|k-1}$, $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ can be computed with
+$\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\!.$ (42)
+Therefore, to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, it is equivalent to reduce the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ since the amount of prior uncertainty $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$ is fixed at time $k$.
 As shown in the next theorem, the amount of uncertainty reduction, i.e., $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, is bounded by the new information extracted from $\mathcal{M}^{x}_{k|k}$.
@@
 The amount of uncertainty reduction at $k$ is
-$\!\!\!\!\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}+\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (38)
+$\!\!\!\!\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}\oplus\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (43)
 satisfying
 $\mathrm{Vol}$ $\left(\Delta\mathcal{Y}_{k|k}\right)\geqslant 2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1},$
-$\!\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant\!\left\|\!A_{3}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
-where $\Delta\mathcal{X}_{k-1|k}=\mathcal{X}_{k-1|k-1}-\mathcal{X}_{k-1|k}$ and $\Delta\mathcal{Y}_{k-1|k}=\mathcal{Y}_{k-1|k-1}-\mathcal{Y}_{k-1|k}$ are the adversary’s uncertainty reduction of $X_{k-1}$ and $Y_{k-1}$, and $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$ quantifies the difference in central estimation with and without considering $\mathcal{X}_{k|k}$, $\left\|A\right\|\frac{\Delta}{=}\sum_{i,j}^{n,m}{\left|a_{i,j}\right|}$ is the absolute value of $A$.
-Proof 3.4.
+$\!\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant$ $\!\left\|\!A_{3}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
+where $\Delta\mathcal{X}_{k-1|k}\!\!=\!\!\mathcal{X}_{k-1|k-1}\!\setminus\!\mathcal{X}_{k-1|k}$, $\Delta\!\mathcal{Y}_{k-1|k}\!\!=\!\!\mathcal{Y}_{k-1|k-1}\!\setminus\!\mathcal{Y}_{k-1|k}$ are the adversary’s uncertainty reduction of $X_{k-1}$ and $Y_{k-1}$, and $\left\|c_{k|k}^{y}\!-\!c_{k|k-1}^{y}\right\|_{1}$ quantifies the difference in central estimation with and without considering $\mathcal{X}_{k|k}$.
+Proof 3.5.
 See Appendix C.
 According to Theorem 9, the reduction of uncertainty $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is highly correlated with the amount of information that the adversary extracts from the observation set $\mathcal{M}_{k|k}^{x}$.
-Furthermore, with Proposition 4 and Theorem 5, we have the following lemma to bound the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k-1|k}\right)$.
+Furthermore, with (42) and Theorem 9, we can show the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k-1|k}\right)$ is bounded in the following lemma.
 Lemma 10.
 The privacy level can be bounded with the following inequalities,
-$\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
+$\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
 $\leqslant\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k}\!\right)\!\leqslant\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$
-Therefore, on the one hand, one can reduce the extracted information $\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$.
-On the other hand, when the private level is high, we also avoid the adversary to update its central estimation $c_{k|k}^{y}$ due to the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$.
+As a result, we can reduce the extracted information $\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ via designing proper observation set $\mathcal{M}^{x}_{k|k}$.
+Moreover, when the privacy level is high, the adversary’s ability to update its central estimate $c_{k|k}^{y}$ is also limited, as indicated by the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$.
+Based on this observation, $\mathcal{M}^{x}_{k|k}$ can also be designed to hinder accurate central estimate updates, further improving the privacy level.
 4 Privacy Filter Design Problem Using the Volumetric Privacy measure
-In this section, we address the privacy filter design problem, where the filter outputs an appropriate observation set that balances the trade-off between the data utility of the public state and the privacy protection of the private state.
+As discussed previously, an inappropriate choice of the observation set would cause privacy leakage of the private state through inference attacks.
+To mitigate this risk, we address the privacy filter design problem in this section.
+The proposed filter determines an appropriate observation set that achieves a desirable balance between preserving the data utility of the public state and ensuring the privacy protection of the private state.
 4.1 The Structure of Privacy Filter
 We begin by defining the decision domain of the privacy filter as follows.
-At time $k$, given the last decision set $\mathcal{X}_{k-1|k-1}$ and the private set $\mathcal{Y}_{k-1|k-1}$, the tightest interval of $X_{k}$ can be computed via Lemma 9, i.e.,
-$\!\!\mathcal{X}_{k|k-1}\!=\!\Psi\!\left(\!A_{1}\!\right)\!\mathcal{X}_{k-1|k-1}\!+\!\Psi\!\left(\!A_{2}\!\right)\!\mathcal{Y}_{k-1|k-1}\!+\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x}\!,$
+At time $k$, given the last decision set $\mathcal{X}_{k-1|k-1}$ and the private set $\mathcal{Y}_{k-1|k-1}$, the inference set of $X_{k}$ can be computed via,
+$\mathcal{Y}_{k|k-1}=A_{3}\mathcal{X}_{k-1|k-1}\oplus A_{4}\mathcal{Y}_{k-1|k-1}\oplus B_{2}\mathcal{W}_{k}^{y},$
 which contains all possible public states that can be reached from any states in $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$.
-Therefore, $\mathcal{X}_{k|k-1}$ is the maximum observation set $\mathcal{M}^{x}_{k|k}$ that the filter can release, i.e., $\mathcal{M}^{x}_{k|k}\subset\mathcal{X}_{k|k-1}$.
+Therefore, $\mathcal{X}_{k|k-1}$ is the maximum observation set $\mathcal{M}^{x}_{k|k}$ that the filter can release, i.e., $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$.
 To maintain the high data utility, the observation set has to satisfy the following constraint,
 $\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}.$
-We next consider the privacy filter drawn in Fig.2, which first randomly generates a set $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$ that contains the actual public state $X_{k}=x_{k}$, and then optimizes the observation set $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ based on linear programming.
-We will demonstrate that the adversary cannot reduce its uncertainty of the public state by reversing the proposed filtering policy.
-Figure 2: The structure of privacy filter.
+To reduce privacy leakage while preserving data utility, we design the privacy filter illustrated in Fig.
+2.
+The design consists of two steps: (1) randomly generate a set $\mathcal{S}_{k|k}^{x}$ such that $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{X}_{k|k-1}$ and $\mathrm{Vol}(\mathcal{S}_{k|k}^{x})\leq\epsilon^{x}$; (2) optimize the observation set $\mathcal{M}^{x}_{k|k}$, which contains $\mathcal{S}_{k|k}^{x}$, to maximize the privacy level.
+Specifically, in the optimization step, given $\mathcal{S}_{k|k}^{x}$, we maximize the privacy level under the inference attack (6)-(12) by solving
+$\mathbf{P_{1}}:$~ $\max_{\mathcal{M}^{x}_{k|k}}~\mathrm{Vol}\!\left(\mathcal{Y}_{k|k}\right)$ (44)
+[…]
+Note that $\mathcal{S}^{x}_{k|k}$ is randomly generated as a subset of $\mathcal{X}_{k|k-1}$.
+In practice, it can be sufficiently small; for instance, it may contain only the true public state $x_{k}$.
+Consequently, recovering $\mathcal{S}^{x}_{k|k}$ could lead to potential privacy leakage and should therefore be avoided.
+In the following, we show that an attacker cannot recover $\mathcal{S}^{x}_{k|k}$ by inverting the optimization problem $\mathbf{P_{1}}$ due to the randomization operation.
+Proposition 11.
+The attacker cannot obtain the smaller set $\mathcal{S}^{x}_{k|k}$ by inverting the optimization problem $\mathbf{P_{1}}$.
+Proof 4.1.
+First, $\mathcal{S}_{k|k}^{x}$ is selected as a random subset of $\mathcal{X}_{k|k-1}$ that contains the true state $x_{k}$.
+Consequently, $x_{k}$ may reside on the boundary of $\mathcal{S}_{k|k}^{x}$.
+Next, let $\mathcal{M}_{k|k}^{x,\star}$ denote the optimal observation set.
+In some cases, $\mathcal{S}_{k|k}^{x}$ may coincide with $\mathcal{M}_{k|k}^{x,\star}$, in which case $x_{k}$ may also lie on the boundary of $\mathcal{M}_{k|k}^{x,\star}$.
+Therefore, an attacker cannot reconstruct a strictly smaller feasible set containing $x_{k}$ by inverting the optimization process.
+As a result, the proposed privacy filter exhibits the following properties: (1) In the absence of inference attacks, the filter output $\mathcal{M}^{x}_{k|k}$ satisfies the utility constraint.
+(2) In the presence of the inference attack described in (6)–(12), the filter output maximizes the privacy level.
+(3) The filter is robust against reverse attacks that attempt to recover the sufficiently small set $\mathcal{S}_{k|k}^{x}$.
+Moreover, since the computational complexity of both the inference attack in (6)–(12) and the volume computation increases with the complexity of the set representations, a trade-off exists between the achievable privacy enhancement of the proposed filter and its computational cost.
+While more sophisticated set representations could improve the accuracy of privacy evaluation, for efficiency and clarity, we next present a concrete design based on the interval approximation described in Sec.
+3.2.
+Figure 2: Structure of the proposed privacy filter.
 4.2 Randomization
@@
 where $\alpha_{k}$ and $\beta_{k}$ are uniform random variables with
-${\alpha}_{k}\in\left[0,\frac{\epsilon^{x}}{2\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}}\right],$
-and
-${\beta}_{k}\in\left[0,\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\right].$
-Since $\left(x_{k}-\underline{X}_{k|k-1}\right)$ is the radius between the actual public state $X_{k}=x_{k}$ and the lower endpoint of $\mathcal{X}_{k|k-1}$, and $\left(\overline{X}_{k|k-1}-x_{k}\right)$ is the radius between $x_{k}$ and the upper endpoint of $\mathcal{X}_{k|k-1}$, the random set ${\mathcal{S}}_{k|k}^{x}$ is a subset of $\mathcal{X}_{k|k-1}$ that contains the actual public state.
-Also, we have $\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ as
-$\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)$
-$=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
+${\alpha}_{k}\!\in\!\left[\!0,\frac{\epsilon^{x}}{2\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}}\right]\!\!,{\beta}_{k}\!\in\!\left[\!0,\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\right]\!\!.$
+Since $\left(x_{k}-\underline{X}_{k|k-1}\right)$ is the radius from the actual public state $X_{k}=x_{k}$ to the lower endpoint of $\mathcal{X}_{k|k-1}$, and $\left(\overline{X}_{k|k-1}-x_{k}\right)$ is the radius from $x_{k}$ to the upper endpoint of $\mathcal{X}_{k|k-1}$, the random set ${\mathcal{S}}_{k|k}^{x}$ becomes a subset of $\mathcal{X}_{k|k-1}$ that contains the actual public state.
+Also, we can shown $\mathcal{S}^{x}_{k|k}$ satisfies the utility constraint as follows,
+$\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
 $\leqslant$ $\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}$
@@
 $=$ $\epsilon^{x}.$
-We next restrict $\mathcal{S}^{x}_{k|k}$ be the subset of the observation set $\mathcal{M}^{x}_{k|k}$, and optimize $\mathcal{M}^{x}_{k|k}$ to enhance the privacy level.
+We next restrict $\mathcal{S}^{x}_{k|k}$ be the subset of the observation set $\mathcal{M}^{x}_{k|k}$, and optimize $\mathcal{M}^{x}_{k|k}$ to improve the privacy level.
 4.3 Privacy Filter Optimization
-[…]
-$\mathbf{P_{1}}:$ $\max_{\mathcal{M}^{x}_{k|k}}\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ (42)
-[…]
-where $\mathcal{S}^{x}_{k|k}\subseteq\mathcal{M}^{x}_{k|k}$ is required by the structure of the privacy filter, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ describes the prior space of $\mathcal{M}^{x}_{k|k}$, and $\mathrm{Vol}\left({\mathcal{M}}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ enforces the requirement of data utility.
-We next demonstrate that the optimization problem $\mathbf{P_{1}}$ can be solved via linear programming.
+In this subsetion, we demonstrate that the optimization problem $\mathbf{P_{1}}$ based on the interval inference can be solved via linear programming.
 Theorem 12.
 The privacy filter optimization problem $\mathbf{P_{1}}$ is equivalent to the following linear programming
-$\mathbf{P_{2}}:\max_{\epsilon^{y},\mathcal{M}^{x}_{k|k},p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}}\epsilon^{y}$ (48)
+$\mathbf{P_{2}}:$ $\max_{\epsilon^{y},\mathcal{M}^{x}_{k|k},p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}}\epsilon^{y}$
 […]
-$\left\{\begin{array}[]{c}p_{k-1|k}^{\Delta z}\geqslant 0\\ p_{k-1|k}^{\Delta z}\geqslant p_{k-1|k-1}^{z}-p_{k-1|k}^{m,z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\\ \end{array}\right.,\text{for}\;Z=X,Y$ (59)
-where $\epsilon^{y}\geqslant 0$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{R}^{2n_{x}}$ and $p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}\in\mathcal{R}^{n_{x}}$.
-Proof 4.1.
+$\left\{\begin{array}[]{c}p_{k-1|k}^{\Delta z}\geqslant 0\\ p_{k-1|k}^{\Delta z}\geqslant p_{k-1|k-1}^{z}-p_{k-1|k}^{m,z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}\\ 2p_{k-1|k}^{\Delta z}\geqslant\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\\ \end{array}\right$., (63)
+where $p_{k-1|k}^{\Delta x}\!\in\mathcal{R}^{n_{x}}$, $Z=X,Y$, $\epsilon^{y}\geqslant 0$ and $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{R}^{2n_{x}}$.
+Proof 4.2.
 See Appendix D
-Figure 3: Inference attack after the optimal privacy filter with $\mathrm{Vol}\left(\mathcal{M}_{k|k}^{x}\right)\leqslant 0.01$.
-Figure 4: Inference attack after the optimal privacy filter with $\mathrm{Vol}\left(\mathcal{M}_{k|k}^{x}\right)\leqslant 0.5$.
-As a result, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 2.3.
-Although the linear programming problem $\mathbf{P_{2}}$ is deterministic, we can demonstrate that the adversary cannot reduce its uncertainty by reversing the optimization process.
-Proposition 8.
-If the adversary infers the private state via the inference attack (2)-(32), then it cannot find a smaller interval subset of $\mathcal{M}_{k|k}^{x,\star}$ containing the actual public state $X_{k}=x_{k}$ via reverting the linear programming.
-Proof 4.2.
-On the one hand, the actual public state $X_{k}=x_{k}$ can be any element of $\mathcal{X}_{k|k-1}$, and $S_{k}^{x}$ is a random subset of $\mathcal{X}_{k|k-1}$ containing $x_{k}$.
-Thus, the upper or lower endpoint of $S_{k}^{x}$ may coincide with $x_{k}$.
-On the other hand, since $\mathcal{M}_{k|k}^{x,\star}$ is the optimal observation set obtained from the linear programming, for any $S_{k}^{x}$ in the vicinity of $\mathcal{M}_{k|k}^{x,\star}$, i.e., $S_{k}^{x}\in\left\{S_{k}^{x}\left|\lim_{\left\|\Delta\right\|_{1}\rightarrow 0}{S_{k}^{x}+\Delta=\mathcal{M}_{k|k}^{x,\star}}\right.\right\}$, the optimal solution remains $\mathcal{M}_{k|k}^{x,\star}$.
-As a result, it is possible that the upper or lower endpoint of $\mathcal{M}_{k|k}^{x,\star}$ corresponds to the actual public state.
-Thus, the minimal interval observation set for the adversary is $\mathcal{M}_{k|k}^{x,\star}$.
-According to Proposition 8, even if the adversary knows the structure of the privacy filter, it cannot revert the linear programming to obtain a smaller interval containing $x_{k}$ to reduce uncertainty.
-In other words, the privacy filer is robust against the reverse optimization attack.
+Figure 3: Inference attack results after applying the optimal privacy filter: (a), (b) Estimated $X_{k}$ and (e), (f) Estimated $Y_{k}$ for $\mathrm{Vol}\!\left(\mathcal{M}_{k|k}^{x}\right)\leq 0.01$; (c), (d) Estimated $X_{k}$ and (g), (h) Estimated $Y_{k}$ for $\mathrm{Vol}\!\left(\mathcal{M}_{k|k}^{x}\right)\leq 0.5$.
+Consequently, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 2.3.
 5 Numerical Verification
@@
 $B_{1}=\left[\begin{matrix}-1.00&0.00\\ 0.00&-1.00\\ \end{matrix}\right],B_{2}=\left[\begin{matrix}4.20&0.00\\ 0.00&2.40\\ \end{matrix}\right],$
-$\left(\mathcal{W}_{k}^{x}\right)^{\top}=\left[\begin{matrix}1.74&1.91&1.94&2.01\\ \end{matrix}\right],$
-$\left(\mathcal{W}_{k}^{y}\right)^{\top}=\left[\begin{matrix}0.91&0.23&0.95&0.43\\ \end{matrix}\right].$
+$\left(\!\mathcal{W}_{k}^{x}\!\right)^{\top}\!\!=\!\left[\!\begin{matrix}1.74&1.91&1.94&2.01\\ \end{matrix}\!\right]\!\!,\left(\!\mathcal{W}_{k}^{y}\!\right)^{\top}\!\!=\!\left[\!\begin{matrix}0.91&0.23&0.95&0.43\\ \end{matrix}\!\right]\!\!.$
 The initial state sets are assumed to be
@@
 In our simulation, the initial states are uniformly sampled from the bounded sets (64)-(65).
-To simulate the approximate periodic fluctuations in demand and productivity, the actual process noises are set to be
+To simulate the approximate periodic fluctuations in demand and productivity, the actual disturbance are set to be
 $\left(W_{k}^{x}\right)^{\top}=\left[\begin{matrix}1.88+0.03\cos\left(\frac{2\pi k}{30+7\rho_{k}}\right)&1.94\\ \end{matrix}\right],$
@@
 As discussed in Section 2, the production rate is private but the inventory information has to be released.
-We first plot the trajectories of system states and their interval tubes in Fig.3 and Fig.4 under the optimal privacy filter design for different values of $\epsilon^{x}$.
+We first plot the trajectories of system states and their interval tubes in Fig.3 under the optimal privacy filter design for different values of $\epsilon^{x}$.
 We also use the central point of the posterior intervals as one of the possible testing estimation, e.g., $\frac{\overline{X}_{k|k}+\underline{X}_{k|k}}{2}$ for $\mathcal{X}_{k|k}$.
@@
 As shown in Fig.3, when $\epsilon^{x}=0.01$ is small, the adversary’s uncertainty about the private production rate is small, and its central estimation closely matches the actual production rate.
-However, when $\epsilon^{x}$ increases to $0.5$, the utility of the inventory information decreases slightly, but this leads to higher uncertainty of the inference attack, causing the adversary’s estimation of the production rate to become less accurate.
+However, when $\epsilon^{x}$ increases to $0.5$, the utility of the inventory information decreases slightly, but this leads to higher uncertainty of the inference attack, causing the adversary’s central estimation of the production rate to become less accurate.
+This observation numerically verifies Theorem 9, confirming that a higher privacy level prevents the adversary from refining its incorrect central estimation.
 Therefore, the proposed privacy filter effectively reduces the privacy leakage of the production rate, though at the cost of introducing some inaccuracies in the inventory information.
 Figure 4: The truncated Gaussian mechanism.
-We analyze the utility-privacy trade-off performance of the optimal privacy filtering policy and compare it with the truncated Gaussian mechanism, which is proposed in [18] for differential privacy.
-In the truncated Gaussian mechanism, illustrated in Fig.
-5, the original state $x_{k}$ is perturbed by additive noise $v_{k}$ drawn from a truncated Gaussian distribution defined over the interval $\left[-\epsilon^{x}/2,\epsilon^{x}/2\right]$, with zero mean and variance $\left(\epsilon^{x}\right)^{2}$.
-Subsequently, the perturbed observation set is released as follows:
-$\mathcal{M}_{k|k}^{x}=\left[\begin{array}[]{c}z_{k}-\epsilon^{x}/\left(2n_{x}\right)\cdot\mathbf{1}_{n_{x}}\\ z_{k}+\epsilon^{x}/\left(2n_{x}\right)\cdot\mathbf{1}_{n_{x}}\\ \end{array}\right].$ (62)
-Since the noise $v_{k}$ is constrained within $\left[-\epsilon^{x}/2,\epsilon^{x}/2\right]$, the released public state $x_{k}$ is guaranteed to reside within the interval $\mathcal{M}_{k|k}^{x}$, and the utility of $x_{k}$ is quantified by $\epsilon^{x}$.
-The privacy-utility trade-off is evaluated by plotting the privacy level of the production rate against the utility of the inventory under different approaches, as depicted in Fig.
-6.
-The results indicate that increasing the utility of inventory leads to a reduction in the privacy level of the production rate, highlighting the inherent trade-off between data utility and privacy protection.
-Compared to the truncated Gaussian mechanism, the optimal privacy filter achieves a higher privacy level while maintaining lower data distortion.
-Figure 6: The privacy level of the private state and utility of the public state.
+We also evaluate the utility–privacy trade-off achieved by the proposed optimal privacy-filtering policy and compare it with two benchmark mechanisms: the noiseless quantization method presented in [21] and the truncated Gaussian mechanism for differential privacy introduced in [31].
+In the quantization-based approach, the state $x_{k}$ is processed through a static quantizer that satisfies the utility constraint, and the quantization bin containing $x_{k}$ is publicly released as $\mathcal{M}_{k|k}^{x}$.
+In contrast, the truncated Gaussian mechanism, illustrated in Fig.
+4, perturbs the original state $x_{k}$ with additive noise $v_{k}$ drawn from a zero-mean truncated Gaussian distribution supported on the interval $\left[-\epsilon^{x}/2,\,\epsilon^{x}/2\right]$ and having variance $\left(\epsilon^{x}\right)^{2}$.
+The perturbed observation set is then released in the form
+$\mathcal{M}_{k|k}^{x}=\left[\begin{array}[]{c}z_{k}-\frac{\epsilon^{x}}{2n_{x}}\cdot\mathbf{1}_{n_{x}}\\ z_{k}+\frac{\epsilon^{x}}{2n_{x}}\cdot\mathbf{1}_{n_{x}}\end{array}\right],$
+where $n_{x}$ denotes the dimension of $x_{k}$ and $\mathbf{1}_{n_{x}}$ is the all-ones vector of length $n_{x}$.
+Because the additive noise $v_{k}$ is bounded within $\left[-\epsilon^{x}/2,\,\epsilon^{x}/2\right]$, the publicly released state is guaranteed to lie within the interval $\mathcal{M}_{k|k}^{x}$.
+The privacy-utility trade-off was evaluated by plotting the average privacy level of the production rate against the average utility of the inventory over 100 random trajectories with horizon 100, as illustrated in Fig.
+5.
+For a fair and clear comparison, the privacy level and utility values for the truncated Gaussian mechanism are normalized to the range $[0,1]$, and its same scaling parameters are applied to the other two mechanisms.
+The results demonstrate that an increase in inventory utility corresponds to a reduction in the privacy level of the production rate, thereby confirming the intrinsic trade-off between data utility and privacy protection.
+Besides, the quantization mechanism and the truncated Gaussian mechanism have similar performance in privacy-utility trade-off.
+Note that these two mechanisms assume static domains of states, while the sets of states in systems subject to disturbance can be estimated and described more precise as addressed by the inference attack.
+The proposed volumetric approach estimates the time-varying private state set and adjusts the output accordingly.
+Furthermore, as discussed in [31], if the volume (e.g., interval length) of domain satisfies certain conditions, the truncated Gaussian mechanism ensures differential privacy.
+However, given the volume constraint, the shape of the public state set can be arbitrary, and certain shapes may lead to substantial volumetric leakage of the private state after inference attacks.
+The proposed volumetric method explicitly accounts for this by considering the geometry of the set based on the assumed inference attack, not only its volume.
+Therefore, it achieves higher privacy levels while maintaining lower data distortion compared with the two static mechanisms.
+It is worth noting that an adversary could employ more sophisticated estimation techniques, e.g., CCG-based approximation, to infer the private set more accurately from the interval observations provided by the privacy filter.
+Nevertheless, as shown in Fig.
+6, the proposed privacy filter still outperforms the other two mechanisms, leveraging knowledge of the underlying state evolution to reduce conservativeness.
+Figure 5: Interval-based inference given the interval privacy filter: the privacy level of the private state and utility of the public state.
+Figure 6: CCG-based inference given the interval privacy filter: the privacy level of the private state and utility of the public state.
 6 Conclusion
-In this paper, we developed a volumetric privacy measure for dynamic systems with UBB noise.
-We defined the inference attack that an adversary uses to estimate the private state and introduced a volumetric measure to evaluate the privacy level.
-We provided computational approaches based on interval analysis and analyzed the theoretical properties of the proposed measure.
-Furthermore, we proposed an optimization-based approach for privacy filter design to defend the system against inference attacks.
-The effectiveness of our method was demonstrated through a production-inventory case study.
+In this paper, we develop a volumetric framework for privacy analysis and defense in dynamic systems subject to bounded disturbance.
+An inference attack, whereby an adversary estimates the private state, is formalized, and a volumetric measure is introduced to quantify the resulting privacy level.
+We develop computational methods based on interval analysis, and establish the theoretical properties of the measure.
+Furthermore, we propose an optimization-based approach for privacy filter design to defend the system against inference attacks.
+The effectiveness of our method is demonstrated through a production-inventory case study.
+It is noted that the performance of the volumetric privacy measure inherently depends on the selected set-membership estimation techniques, and its evaluation accuracy varies with different set representations.
+Future research will focus on developing approximation methods that ensure improved accuracy, robustness and broader applicability.
 Appendix A Proof of Lemma 6
 At the time step $k=0$, the adversary only has prior knowledge, i.e., $\mathcal{Y}_{0|-1}$, therefore, its inference set is $\mathcal{Y}_{0|0}=\mathcal{Y}_{0|-1}$.
+Also, since at the time step $k=0$, the backward calibration (6) and (7) is not available, the adversary can only calibrate the public state set with its prior knowledge $\mathcal{X}_{0|-1}$ and the observation set $\mathcal{M}_{0|0}^{x}$ according to (40).
 To prove Lemma 6 for $k\geqslant 1$, we need the following lemma that computes the tightest interval by forward reachability analysis.
@@
 [32, 33] Consider the static system $S=AM+BW$, where $M$ and $W$ are bounded intervals, the tightest interval for $S$, i.e., $\mathcal{S}$ can be computed as
-$\mathcal{S}=\Psi\left(A\right)\mathcal{M}+\Psi\left(B\right)\mathcal{W},$ (63)
-where
-$\Psi\left(N\right)=\left[\begin{matrix}\frac{N+\left|N\right|}{2}&\frac{N-\left|N\right|}{2}\\ \frac{N-\left|N\right|}{2}&\frac{N+\left|N\right|}{2}\\ \end{matrix}\right].$ (64)
-Also, we can compute the radius and center of the intervl $\mathcal{S}$ via
-$p^{s}=\left|A\right|p^{m}+\left|B\right|p^{w},$ (65)
-$c^{s}=Ac^{m}+Bc^{w}.$ (66)
-When $A_{1}$ and $A_{2}$ are invertible, we have
-$X_{k-1}=A_{1}^{-1}X_{k}+\left(-A_{1}^{-1}A_{2}\right)Y_{k-1}+\left(-A_{1}^{-1}B_{1}\right)W_{k}^{x},$
-and
-$Y_{k-1}=A_{2}^{-1}X_{k}+\left(-A_{2}^{-1}A_{1}\right)X_{k-1}+\left(-A_{2}^{-1}B_{1}\right)W_{k}^{x}.$
-According to Lemma 9, the tightest intervals for (2.3) and (2.3) are (2) and (2).
+$\mathcal{S}=\Psi\left(A\right)\mathcal{M}\oplus\Psi\left(B\right)\mathcal{W}.$
+Also, we can compute its radius and center via
+$p^{s}=\left|A\right|p^{m}$ $+\left|B\right|p^{w},$
+$c^{s}=Ac^{m}$ $+Bc^{w}.$
+According to Lemma 13, the tightest intervals for (6) and (7) are (24) and (25).
 Then, the intersection of different intervals, i.e., (8) and (9), can be computed with (28) and (31).
 Finally, the one-step forward reachable set (10) can be approximated with the tightest interval (36) based on Lemma 13, and the calibrated uncertainty set $\mathcal{X}_{k|k}$ and the tightest prior inference set $\mathcal{Y}_{k|k-1}$ can be approximated similarly.
-Finally, since at the time step $k=0$, the backward calibration (2.3) and (2.3) is not available, the adversary can only calibrate the public state set with its prior knowledge $\mathcal{X}_{0|-1}$ and the observation set $\mathcal{M}_{0|0}^{x}$ according to (36).
 Appendix B Proof of Lemma 8
 According to Lemma 13, the radius of $\mathcal{M}^{y}_{k-1|k}$ can be computed as
-$p_{k-1|k}^{m,y}=\left|A_{2}^{-1}\right|p_{k|k}^{m,x}+\left|A_{2}^{-1}A_{1}\right|p_{k-1|k-1}^{x}+\left|A_{2}^{-1}B_{1}\right|p_{k}^{w,x},$
+$p_{k-1|k}^{m,y}\!=\!\left|\!A_{2}^{-1}\!\right|p_{k|k}^{m,x}\!+\!\left|\!A_{2}^{\!-\!1}\!A_{1}\!\right|p_{k\!-\!1|k-\!1}^{x}\!+\!\left|\!A_{2}^{\!-\!1}\!B_{1}\!\right|p_{k}^{w,x}\!,\!\!$ (66)
 where $p^{m,x}_{k|k}$, $p_{k-1|k-1}^{x}$ and $p^{w,x}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{X}_{k-1|k-1}$ and $\mathcal{W}_{k}^{x}$, respectively.
@@
 Also, the radius of $\mathcal{Y}_{k|k}$ can be computed as
-$p_{k|k}^{y}=\left|A_{3}\right|p_{k-1|k}^{x}+\left|A_{4}\right|p_{k-1|k}^{y}+\left|B_{2}\right|p_{k}^{w,y}.$
-Therefore, with (B), we have
+$p_{k|k}^{y}=\left|A_{3}\right|p_{k-1|k}^{x}+\left|A_{4}\right|p_{k-1|k}^{y}+\left|B_{2}\right|p_{k}^{w,y}.$ (67)
+By substituting (66) and $p_{k-1|k}^{y}\leqslant p_{k-1|k}^{m,y}$ into (67), we have
 $p_{k|k}^{y}\leqslant\left|A_{3}\right|p_{k-1|k}^{x}+\left|B_{2}\right|p_{k}^{w,y}+\left|A_{4}\right|\left|A_{2}^{-1}B_{1}\right|p_{k}^{w,x}$
@@
 The difference set $\Delta\mathcal{Y}_{k|k}$ is computed as,
-$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}=\Phi\left(A_{3}\right)\Delta\mathcal{X}_{k-1|k}+\Phi\left(A_{4}\right)\Delta\mathcal{Y}_{k-1|k},$
+$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\setminus\mathcal{Y}_{k|k}=\Phi\left(A_{3}\right)\Delta\mathcal{X}_{k-1|k}\oplus\Phi\left(A_{4}\right)\Delta\mathcal{Y}_{k-1|k},$
 where
-$\Delta\mathcal{X}_{k-1|k}$ $=\mathcal{X}_{k-1|k-1}-\mathcal{M}_{k-1|k}^{x}$
+$\Delta\mathcal{X}_{k-1|k}$ $=\mathcal{X}_{k-1|k-1}\setminus\mathcal{X}_{k-1|k}$
 $=\left[\begin{array}[]{c}\min\left\{\underline{X}_{k-1|k-1}-\underline{M}_{k-1|k}^{x},0\right\}\\ \max\left\{\overline{X}_{k-1|k-1}-\overline{M}_{k-1|k}^{x},0\right\}\\ \end{array}\right],$ (70)
-$\Delta\mathcal{Y}_{k-1|k}$ $=\mathcal{Y}_{k-1|k-1}-\mathcal{Y}_{k-1|k}$
+$\Delta\mathcal{Y}_{k-1|k}$ $=\mathcal{Y}_{k-1|k-1}\setminus\mathcal{Y}_{k-1|k}$
 $=\left[\begin{array}[]{c}\min\left\{\underline{Y}_{k-1|k-1}-\underline{M}_{k-1|k}^{y},0\right\}\\ \max\left\{\overline{Y}_{k-1|k-1}-\overline{M}_{k-1|k}^{y},0\right\}\\ \end{array}\right].$ (73)
@@
 With Lemma 13, we have
-$p_{k|k}^{\Delta y}=\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y},$ (73)
+$p_{k|k}^{\Delta y}=\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y},$
 where the radius $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ can be computed via
@@
 $=$ $\max\left\{\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},0\right\}-\min\left\{\underline{Z}_{k-1|k-1}-\underline{M}_{k-1|k}^{z},0\right\}$
-$=$ $\max\left\{\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},0\right\}+\max\left\{\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1},0\right\}$
 $=$ $\max\left\{0,\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z}+\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1},\right$.
-$\left.\overline{Z}_{k-1|k-1}-\overline{M}_{k-1|k}^{z},\underline{M}_{k-1|k}^{z}-\underline{Z}_{k-1|k-1}\right\},\text{for}\,Z=X,Y,$ (74)
+$\!\!\!\!\overline{Z}_{k-\!1|k-\!1}\!-\!\overline{M}_{k-\!1|k}^{z},\underline{M}_{k-1|k}^{z}\!-\!\underline{Z}_{k-\!1|k-\!1}\left.\right\},\text{for}\,Z=X,Y,\!\!$ (74)
 which satisfies $p_{k-1|k}^{\Delta z}\geqslant 0$. As a result, we have
-$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$
-$=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
-=$\overset{(a)}{=}\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
-$\leqslant\left\|A_{3}\right\|\left\|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|A_{4}\right\|\left\|p_{k-1|k}^{\Delta y}\right\|_{1}$
-$=\left\|A_{3}\right\|\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right),$ (75)
+$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ $=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
+$\overset{(a)}{\leqslant}\left\|A_{3}\right\|_{1}\left\|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|A_{4}\right\|_{1}\left\|p_{k-1|k}^{\Delta y}\right\|_{1}$
+$=\left\|A_{3}\right\|_{1}\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|_{1}\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right),$
 where $(a)$ is due to $p_{k-1|k}^{\Delta x}\geqslant 0$ and $p_{k-1|k}^{\Delta y}\geqslant 0$.
 Besides, given an interval $\mathcal{X}$, we can express it with its center point and radius, i.e., $\mathcal{X}=\left[\begin{array}[]{c}\frac{c-p}{2}\\ \frac{c+p}{2}\end{array}\right]$.
-Therefore, we have (C),
-$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$
-$=$ $\left\|\overline{Y}_{k|k-1}-\overline{Y}_{k|k}\right\|_{1}+\left\|\underline{Y}_{k|k-1}-\underline{Y}_{k|k}\right\|_{1}$
-$=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
+Therefore, we have
+$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
 $\geqslant$ $\left\|\overline{Y}_{k|k}+\underline{Y}_{k|k}-\left(\overline{Y}_{k|k-1}+\underline{Y}_{k|k-1}\right)\right\|_{1}$
-$\geqslant$ $2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$ (76)
+$\geqslant$ $2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$
 Appendix D Proof of Theorem 12

### v5

@@
 Abstract
-This paper presents a volumetric privacy framework for dynamical systems subject to bounded disturbances, developed without requiring prior knowledge of their probability distributions.
-We consider systems with both public and private states, where a set containing the public state is shared as the observation.
-An adversary is assumed to execute an inference attack by exploiting the observed public state set to estimate an uncertainty set for the private state.
-The volume of this inferred set quantifies the adversary’s estimation uncertainty and serves as the proposed volumetric privacy metric.
-Approximate set-membership estimation techniques are developed to compute the private-state uncertainty set, and the properties of the privacy measure are analyzed, demonstrating that it is bounded by the information gain from the observation set.
-Furthermore, an optimization-based privacy filter design problem is formulated, employing randomization and linear programming to enhance the volumetric privacy level.
+In this paper, we first present a volumetric privacy measure for dynamical systems with bounded disturbances, wherein the states of the system contain private information and an adversary with access to sensor measurements attempts to infer the set of potential values of the private information.
+Under the proposed privacy measure, the volume of the uncertainty set of the adversary given the sensor measurements is considered as the privacy level of the system.
+We next characteristic the time evolution of the proposed privacy measure and study its properties for a particular system with both public and private states, where a set containing the public state is shared as the observation.
+Approximate set-membership estimation techniques are developed to compute the private-state uncertainty set, and the properties of the privacy measure are analyzed, demonstrating that the uncertainty reduction of the adversary is bounded by the information gain from the observation set.
+Furthermore, an optimization-based privacy filter design problem is formulated, employing randomization and linear programming to enhance the privacy level.
 The effectiveness of the proposed approach is validated through a production–inventory case study.
@@
 Data sharing plays a pivotal role in enabling cooperative decision-making and optimization in dynamic processes.
-However, the exposure of such data may inadvertently reveal sensitive information.
-Specifically, correlations between shared metrics and underlying operational information can be exploited by adversaries to develop competitive and malicious strageties.
-This challenge highlights the critical need for methodologies that preserve data utility while ensuring privacy protection for dynamic systems.
-Dynamical systems subject to bounded disturbances without knowledge of their underlying distributions provide a natural framework for modeling numerous practical applications involving sensitive information, e.g., inventory–production systems and connected vehicles.
-However, the notion of privacy in such systems remains insufficiently explored.
-Within the context of set-membership estimation, the states of these systems can be represented by geometric sets, such as ellipsoids or zonotopes, whose volumes quantify the degree of inference uncertainty.
+However, the exposure of such data may inadvertently reveal sensitive information[1].
+Dynamical systems subject to bounded disturbances without knowledge of their underlying distributions provide a natural framework for modeling numerous practical applications involving sensitive information.
+Despite their importance, the notion of privacy in such systems remains insufficiently explored.
+Within the context of set-membership estimation, the states of systems with bounded disturbance can be represented by geometric sets, such as ellipsoids or zonotopes, whose volumes quantify the degree of inference uncertainty.
 Motivated by these considerations, this paper investigates the notion of volumetric privacy for systems affected by bounded disturbance.
@@
 As discussed in Sec.
-3.1, the complexity of CCG-based inference grows exponentially with time, motivating our use of interval analysis for computational efficiency.
+3.2, the complexity of CCG-based inference grows exponentially with time, motivating our use of interval analysis for computational efficiency.
 Noiseless privacy [21] and non-stochastic privacy [23] employed non-stochastic information-theoretic approaches to limit information leakage, assuming static private-variable domains and without accounting for temporal dependencies in sequential data.
-While noiseless privacy, non-stochastic privacy, and our volumetric privacy all achieve privacy through the release of bounded outputs.
-However, in our setup, dynamical systems subject to bounded disturbance have time-variant private-state reachable sets that can be recursively estimated, enabling dynamic leakage evaluation.
+While noiseless privacy, non-stochastic privacy, and our volumetric privacy all achieve privacy through the release of bounded outputs, in our setup, dynamical systems subject to bounded disturbance have time-variant private-state reachable sets that can be recursively estimated, enabling dynamic leakage evaluation.
 Building on this insight, the proposed volumetric privacy filter dynamically evaluates the private state set and adapts the observation accordingly, thereby achieving higher privacy levels with lower data distortion, as shown in Sec.
@@
 For example, the authors in [24] established a connection between network privacy and its observability space, proposing a privacy-aware communication protocol that achieves average consensus while protecting initial states.
-In [3], the trace of the observability Gramian was employed to quantify information leakage through an intruder node, and an online optimization strategy was proposed to adapt communications in order to minimize such leakage.
-Privacy-preserving consensus designs were also proposed in [25, 26], including augmented states and novel consensus algorithms with simultaneous accuracy, resilience, and privacy guarantees.
+The authors in [25, 26] presented privacy-preserving consensus algorithms that incorporate augmented states and novel consensus mechanisms, balancing accuracy, resilience, and privacy guarantees simultaneously.
 State decomposition methods, as in [27], split each node’s state into randomized components to prevent disclosure of individual states during consensus.
 1.3 Contributions
-This paper studies privacy for dynamic systems with bounded disturbance as shown in Fig.
-1, where the system state is split into public $X_{k}$ and private $Y_{k}$, and an adversary uses the public observation set $\mathcal{M}_{k|k}^{x}$ to infer $Y_{k}$ via an uncertainty set $\mathcal{Y}_{k|k}$.
+In this paper, we investigate volumetric privacy in dynamical systems subject to bounded disturbances, as illustrated in Fig.
+1.
+The system state $S_{k}$ contains both utility information $X_{k}$ and private information $Y_{k}$, while an adversary exploits the observation set $\mathcal{M}_{k|k}^{x}$ to infer the private information $Y_{k}$ via the associated uncertainty set $\mathcal{Y}_{k|k}$.
 Figure 1: The system setup.
@@
 1.4 Outline
-The rest of the paper is organized as follows.
-Section 2 introduces the system model, the inference attack and defines the volumetric privacy and utility measure.
-Section 3 provides computational approaches for privacy level evaluation, and discusses the properties of the proposed measure.
-Section 4 presents an optimal privacy filter design to mitigate privacy leakage while maintaining a certain utility level.
-Section 5 presents numerical results, followed by the conclusions in Section 6.
+The remainder of the paper is organized as follows.
+Section 2 introduces the system model and defines the volumetric privacy and utility measures.
+Section 3 formalizes the inference attack, presents computational approaches for evaluating the privacy level, and discusses the properties of the proposed measure.
+Section 4 develops an optimal privacy filter to mitigate privacy leakage while preserving a desired utility level.
+Numerical results are presented in Section 5, followed by concluding remarks in Section 6.
 1.5 Notation
@@
 The vector $\mathbf{1}_{n_{x}}$ denotes a column vector of ones with $n_{x}$ dimensions, while $I_{n_{x}\!\times\!n_{x}}$ represents an identity matrix of size $n_{x}\times n_{x}$.
-The operator $diag(v)$ denotes a diagonal matrix constructed from the vector $v$.
-2 System Model and Inference Attack
+The operator $\mathrm{diag}(v)$ denotes a diagonal matrix constructed from the vector $v$.
+2 System Model and Privacy Measure
 2.1 System model
-We consider the following stable system model $\mathbf{G_{1}}$,
-$\mathbf{G_{1}}:\left\{\begin{array}[]{c}X_{k}=A_{1}X_{k-1}+A_{2}Y_{k-1}+B_{1}W_{k}^{x}\\ Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+B_{2}W_{k}^{y}\\ \end{array}\right$., (3)
-where $A_{1}$ and $A_{2}$ are invertible, $Y_{k}\in\mathcal{R}^{n}$ is the private state, $X_{k}\in\mathcal{R}^{n}$ is the public state to be released, $W_{k}^{x}\in\mathcal{W}_{k}^{x}\subseteq\mathcal{R}^{m}$ and $W_{k}^{y}\in\mathcal{W}_{k}^{y}\subseteq\mathcal{R}^{m}$ are the unknown disturbance with bounded sets $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$.
-Note that the underlying probability distributions of $W_{k}^{x}$ and $W_{k}^{y}$ are unknown.
-Besides, the initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively.
-We assume that the adversary has full knowledge of system model $\mathbf{G_{1}}$ and will collect information of the public state to infer the private state.
+Consider the following stable dynamical system
+$\mathbf{G}:\left\{\begin{array}[]{l}S_{k}=f\left(S_{k-1},W_{k}\right),\\[3.0pt] O_{k}=g\left(S_{k},V_{k}\right),\end{array}\right$. (3)
+where $S_{k}$ denotes the system state, with initial condition $S_{0}$ belonging to a bounded set $\mathcal{S}_{0}\subset\mathcal{R}^{s}$, $O_{k}$ is the sensor measurement.
+The unknown disturbances $W_{k}$ and $V_{k}$ are assumed to be bounded within the sets $\mathcal{W}_{k}\subset\mathcal{R}^{s}$ and $\mathcal{V}_{k}\subset\mathcal{R}^{o}$, respectively.
+The system state $S_{k}$ contains both private and utility-related information, which are represented as continuous variables
+$Y_{k}=h(S_{k}),\qquad X_{k}=u(S_{k}),$ (4)
+where $Y_{k}$ represents the private information that must be kept confidential, and $X_{k}$ corresponds to the utility information intended to be disclosed.
+We assume that the adversary has full knowledge of system model $\mathbf{G}$ and will collect observations of the system to infer the private information.
+A special case of this model is the following linear non-Gaussian system with invertible $A_{1}$ and $A_{2}$,
+$\mathbf{G}_{1}:\left\{\begin{array}[]{c}X_{k}=A_{1}X_{k-1}+A_{2}Y_{k-1}+B_{1}W_{k}^{x}\\ Y_{k}=A_{3}X_{k-1}+A_{4}Y_{k-1}+B_{2}W_{k}^{y}\\ O_{k}=X_{k}\\ \end{array}\right$., (8)
+where the system state $S_{k}=[X_{k}^{\top},\,Y_{k}^{\top}]^{\top}$ consists of the public state $X_{k}\in\mathcal{R}^{n_{x}}$ and the private state $Y_{k}\in\mathcal{R}^{n_{x}}$, and the measurement $O_{k}$ corresponds to the public state $X_{k}$.
+The initial public and private states belong to $\mathcal{X}_{0|-1}$ and $\mathcal{Y}_{0|-1}$, respectively.
 2.2 Motivating Examples
-In this subsection, we consider two motivating examples to illustrate the necessity of protecting privacy of systems $\mathbf{G_{1}}$.
+We next consider two motivating examples to illustrate the necessity of protecting privacy of systems with bounded disturbance.
 Production-inventory system: In supply chain management [28, 29], the inventory level $X_{k}$ and the production rate $Y_{k}$ evolves according to $\mathbf{G}_{1}$.
@@
 To protect location privacy, vehicles may intentionally report blurred or randomized velocity observations that preserve system utility but reduce the risk of precise location inference.
-2.3 Inference Attack
-Due to the presence of unknown disturbance terms belonging to bounded sets, multiple private states may correspond to the same public state.
-Consequently, for a given public state set, an adversary cannot determine the exact private state; instead, it can only identify a corresponding uncertainty set.
-we next define how the adversary infers the private state using set-theoretic operations.
-We assume that the adversary observes a set of public states, denoted as $\mathcal{M}_{k|k}^{x}$, which includes the actual public state value $X_{k}=x_{k}$ and other elements to obfuscate the adversary’s estimation of private states.
-Based on the observed public state set, the adversary identifies all potential values of the private state that align with $\mathcal{M}_{k|k}^{x}$ to construct its uncertainty set.
-This process is referred to as the inference attack.
-We next define the inference attack recursively.
+2.3 Privacy and Utility Measures
+Notably, in differential privacy for dynamical systems, the ranges of utility and private information are typically assumed to be time-invariant and unbounded, and privacy is characterized by probabilistic indistinguishability.
+However, for systems subject to bounded disturbances, the utility and private information can be represented by bounded uncertainty sets that are updated online using set-membership estimation.
+Accordingly, it is desirable to introduce a privacy measure tailored to such set-based descriptions.
+Given the sensor measurements, an adversary can employ set-membership estimation techniques, e.g., [18, 19], to estimate both the utility and private information by constructing uncertainty sets $\mathcal{X}_{k|k}$ and $\mathcal{Y}_{k|k}$ that contain the true values $X_{k}=x_{k}$ and $Y_{k}=y_{k}$, respectively.
+We would further provide numerical approaches to illustrate how the adversary estimates $\mathcal{X}_{k|k}$ and $\mathcal{Y}_{k|k}$ in next section.
+This inference procedure, referred to as an inference attack, results in privacy leakage, as the adversary’s confidence in the private information increases when the uncertainty set $\mathcal{Y}_{k|k}$ shrinks.
+In the extreme case where $\mathcal{Y}_{k|k}=\{y_{k}\}$, the private value is fully revealed.
+Since $Y_{k}$ resides in a continuous space, the cardinality of $\mathcal{Y}_{k|k}$ is not meaningful, motivating a geometric approach to quantify uncertainty.
+We therefore employ the volume of the uncertainty set as a quantitative measure of privacy:
+$\mathrm{Vol}\!\left(\mathcal{Y}_{k|k}\right)=\int_{\mathcal{Y}_{k|k}}\mathrm{d}y,$ (9)
+where $\mathrm{Vol}(\mathcal{Y}_{k|k})$ denotes the Lebesgue measure of the set $\mathcal{Y}_{k|k}\subseteq\mathcal{R}^{n_{y}}$. The privacy measure at time $k$ is then defined as
+$P_{k}\!\left(\mathcal{Y}_{k|k}\right):=\mathrm{Vol}\!\left(\mathcal{Y}_{k|k}\right).$ (10)
+A smaller volume corresponds to reduced privacy, while $\mathrm{Vol}(\mathcal{Y}_{k|k}\!)\!=\!0$ indicates complete exposure.
+Similarly, the utility information can be represented by the uncertainty set $\mathcal{X}_{k|k}$.
+Larger $\mathcal{X}_{k|k}$ implies greater uncertainty in recovering the true public information $X_{k}=x_{k}$, resulting in higher distortion.
+Accordingly, we define the utility measure as
+$U_{k}\!\left(\mathcal{X}_{k|k}\right):=\frac{1}{\mathrm{Vol}\!\left(\mathcal{X}_{k|k}\right)},$ (11)
+so that higher $U_{k}$ corresponds to better utility. Throughout, we assume that all relevant uncertainty sets are measurable with finite, positive volume.
+In the following, we analyze the proposed volumetric privacy measure in the context of the linear non-Gaussian system $\mathbf{G}_{1}$, and we develop practical approximations for evaluating the resulting privacy leakage.
+Based on this analysis, we design a privacy filter that mitigates leakage while satisfying utility requirements.
+A extension study of volumetric privacy for general nonlinear systems is left for future work.
+3 Inference Attack for Linear Systems
+3.1 Inference Attack via Set Operations
+We assume that the adversary observes a set of public states, denoted by $\mathcal{M}_{k|k}^{x}$, which is generated by the privacy filter, as shown in Fig.
+1.
+The observation set contains the true public state $X_{k}=x_{k}$ along with additional elements intended to obfuscate the adversary’s estimation of the private states.
+The adversary then performs the inference attack by identifying all possible values of the private state that are consistent with $\mathcal{M}_{k|k}^{x}$, thereby constructing its uncertainty set.
+In the following, we define the inference attack recursively.
 At time $k$, given the public state set $\mathcal{X}_{k-1|k-1}$ and the uncertainty private state set $\mathcal{Y}_{k-1|k-1}$, the set of states can be predicted based on the system model (8), i.e.,
@@
 Moreover, given uncertainty sets $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$ that contain the true system states $X_{k-1}\!=\!x_{k-1}$ and $Y_{k-1}\!=\!y_{k-1}$, if the observation set $\mathcal{M}^{x}_{k|k}$ contains the true public state $X_{k}\!=\!x_{k}$, then the inference set $\mathcal{Y}_{k|k}$ contains the true private state $Y_{k}\!=\!y_{k}$.
-Proof 2.1.
+Proof 3.1.
 Since $\mathcal{M}^{x}_{k|k}$ contains $x_{k}$, it follows from (14) that $x_{k-1}\!\in\!\mathcal{M}^{x}_{k-1|k}$.
@@
 In particular, if the uncertainty set $\mathcal{Y}_{k|k}$ contains only one element, then the adversary can obtain the actual private state.
-2.4 Privacy and Utility Measures
-As discussed in Section 2.3, the uncertainty set of private state $\mathcal{Y}_{k|k}$ encompasses all possible elements that correspond to the same observation set, $\mathcal{M}_{k|k}^{x}$.
-The attacker has more inference uncertainty about the private state if $\mathcal{Y}_{k|k}$ contains more elements.
-However, since the state space is continuous, the number of elements in such sets is uncountable.
-To address this, we propose to use the volume of the uncertainty set as a quantitative measure of privacy.
-Specifically, we define
-$\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)=\int_{\mathcal{Y}_{k|k}}\mathrm{d}y,$ (13)
-where $\mathrm{Vol}(\mathcal{Y}_{k|k})$ denotes the Lebesgue measure (i.e., the geometric volume) of the set $\mathcal{Y}_{k|k}\subseteq\mathcal{R}^{n}$.
-We then define the privacy measure at time $k$ as
-$P_{k}\left(\mathcal{Y}_{k|k}\right):=\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right).$ (14)
-Since the uncertainty set $\mathcal{Y}_{k|k}$ contains the actual value of private state $Y_{k}=y_{k}$, the adversary can accurately access to $y_{k}$ if the volume of $\mathcal{Y}_{k|k}$ is zero.
-To protect the system from inference attack, we can increase the privacy level of the private state via maximizing $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$.
-On the other hand, it causes more utility distortion of the public state if the observation set $\mathcal{M}^{x}_{k|k}$ contains more elements, since it is more difficult for the receiver to recover the actual public state.
-To address this, we would reduce distortion via maximizing the following defined utility of public state set $\mathcal{M}^{x}_{k|k}$,
-$U_{k}\left(\mathcal{M}^{x}_{k|k}\right)=1/\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right).$ (15)
-Notably, set operations in the proposed inference attack and the volume computation are generally computational costly due to the continuous nature of the state space.
-Consequently, it is essential to provide efficient computational tools for privacy level evaluation.
-Also, from a defense perspective, another critical task is the design of effective mechanisms to mitigate privacy leakage.
-To address these challenges, we study computation approaches for inference attack in Sec.
-3, and present an optimal privacy filter design to reduce privacy leakage in Sec.
-4.
-3 Inference Attack Approximation
 As addressed in existing set-membership estimation approaches [18, 19], the set operations involved in inference attacks can be computationally expensive.
@@
 In this subsection, we show that the inference attack can be approximated using the constrained convex generator (CCG), a general set representation proposed in [30].
-Definition 2.
+Definition 2 (CCG Representation).
 [30] The constrained convex generator $\mathcal{Z}=\left(G,c,A,b,\mathcal{C}\right)\subset\mathcal{R}^{n}$ is defined as
-$\mathcal{Z}=\left\{G\xi+c:A\xi=b,\xi\in\mathcal{C}\right\},$ (16)
+$\mathcal{Z}=\left\{G\xi+c:A\xi=b,\xi\in\mathcal{C}\right\},$
 where $G\in\mathcal{R}^{n\times n_{g}}$, $c\in\mathcal{R}^{n}$, $A\in\mathcal{R}^{n_{c}\times n_{g}}$, $b\in\mathcal{R}^{n_{c}}$ and $\mathcal{C}=\left\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{n_{p}}\right\}$, and $\mathcal{C}_{i}\subset\mathcal{R}^{m_{i}}$ are convex sets with $\sum_{i=1}^{n_{p}}{m_{i}}=n_{g}$.
@@
 The computational complexity of the CCG-based inference attack at time $k$ is at least $\mathcal{O}\!\left(c^{k-1}n^{3}\right)$ for some constant $c>1$, and both the column dimension of the generator matrices $G$ and the number of constraints grow exponentially with $k$.
-Proof 3.1.
+Proof 3.2.
 The dominant operation in the inference attack from (14) to (20) is the multiplication of an $n\times n$ matrix with an $n\times m$ matrix, which has computational complexity $\mathcal{O}(mn^{2})$.
@@
 We next consider an interval-based approximation approach for computing the inference sets.
-This approach is a special case of CCG-based inference but significantly reduces both computational and analytical complexity due to the efficiency of interval arithmetic.
-Moreover, interval-based inference serves as the foundation for designing the optimal privacy filter, as discussed in Sec.
-4.
-Definition 5.
-The interval $\mathcal{X}=\{X\mid\underline{X}\leqslant X\leqslant\overline{X}\}$ is defined as $\begin{bmatrix}\underline{X}\\[3.00003pt] \overline{X}\end{bmatrix}$ with the lower and upper bounds $\underline{X}$ and $\overline{X}$.
-An interval $\mathcal{X}$ can also be expressed as a special case of the CCG representation $\mathcal{X}=\left\{\mathrm{diag}(p^{x})\,\xi+c^{x}\;:\;\xi\in\mathcal{R}^{n_{x}},\;\|\xi\|_{\infty}\leqslant 1\right\}$, with the center point $c^{x}=\frac{\overline{X}+\underline{X}}{2}$ and the radius $p^{x}=\frac{\overline{X}-\underline{X}}{2}$.
-Also, the volume of the interval $\mathcal{X}$ is computed as $\mathrm{Vol}\left(\mathcal{X}\right)=\sum_{i=1}^{n}{\left(\overline{X}\left(i\right)-\underline{X}\left(i\right)\right)}$ where $\overline{X}(i)$ and $\underline{X}(i)$ denote the upper and lower bounds of the $i$-th dimension of the interval $\mathcal{X}$, respectively.
-Given a block matrix $A=[A_{1},\,A_{2}]$, the multiplication of $A$ with an interval $\mathcal{X}$ is defined as $A\mathcal{X}=A_{1}\underline{X}+A_{2}\overline{X}$.
-If $\mathcal{X}$ and $\mathcal{Y}$ are intervals, their Minkowski sum is $\mathcal{X}\oplus\mathcal{Y}=\begin{bmatrix}\underline{X}+\underline{Y}\\[3.00003pt] \overline{X}+\overline{Y}\end{bmatrix}$, and their difference, used only for volume evaluation, is $\mathcal{X}\setminus\mathcal{Y}=\begin{bmatrix}\underline{X}-\underline{Y}\\[3.00003pt] \overline{X}-\overline{Y}\end{bmatrix}$.
-The intersection of $\mathcal{X}$ and $\mathcal{Y}$ is $\mathcal{X}\cap\mathcal{Y}=\begin{bmatrix}\max\{\underline{X},\underline{Y}\}\\[3.00003pt] \min\{\overline{X},\overline{Y}\}\end{bmatrix}$.
+This approach can be viewed as a special case of the CCG-based inference, but it significantly reduces both computational and analytical complexity due to the efficiency of interval arithmetic.
+Definition 5 (Interval Representation).
+An interval $\mathcal{X}=\{X\mid\underline{X}\leq X\leq\overline{X}\}$ is equivalently represented as $\mathcal{X}=\left[\begin{matrix}\underline{X}^{\top}&\overline{X}^{\top}\\ \end{matrix}\right]^{\top}$, where $\underline{X}$ and $\overline{X}$ are the lower and upper bounds, respectively.
+Equivalently, an interval can be expressed as a special case of the CCG representation,
+$\mathcal{X}=\left\{\mathrm{diag}(p^{x})\,\xi+c^{x}\;:\;\xi\in\mathbb{R}^{n_{x}},\;\|\xi\|_{\infty}\leq 1\right\},$
+with center $c^{x}=(\overline{X}+\underline{X})/2$ and radius $p^{x}=(\overline{X}-\underline{X})/2$.
+The volume of $\mathcal{X}$ is given by $\mathrm{Vol}(\mathcal{X})=\prod_{i=1}^{n_{x}}\big(\overline{X}(i)-\underline{X}(i)\big)=\prod_{i=1}^{n_{x}}2p^{x}(i)$, where $\overline{X}(i)$ and $\underline{X}(i)$ denote the upper and lower bounds of the $i$-th dimension, respectively.
+To reduce computational complexity, a surrogate measure for the size of $\mathcal{X}$ can be defined as the total length across all dimensions $\overline{\mathrm{Vol}}(\mathcal{X})=\sum_{i=1}^{n_{x}}\big(\overline{X}(i)-\underline{X}(i)\big)=\sum_{i=1}^{n_{x}}2p^{x}(i)$, which is linear in the upper and lower bounds and therefore easier to compute.
+By the inequality of arithmetic and geometric means, this surrogate measure bounds the geometric volume, i.e., $\overline{\mathrm{Vol}}(\mathcal{X})/n\geq\sqrt[n]{\mathrm{Vol}(\mathcal{X})}$.
+Consequently, maintaining a large $\mathcal{X}$ implies a correspondingly large $\overline{\mathrm{Vol}}(\mathcal{X})$, making it a suitable surrogate metric.
+[…]
 We now assume that the uncertainty sets $\mathcal{W}^{x}_{k}$, $\mathcal{W}^{y}_{k}$, $\mathcal{X}_{0|-1}$, and $\mathcal{Y}_{0|-1}$ are represented as intervals.
@@
 $\mathcal{X}_{0|0}=\left[\begin{array}[]{c}\max\left\{\underline{M}_{0|0}^{x},\underline{X}_{0|-1}\right\}\\ \min\left\{\overline{M}_{0|0}^{x},\overline{X}_{0|-1}\right\}\\ \end{array}\right].$ (44)
-Proof 3.2.
+Proof 3.3.
 See Appendix A.
@@
 The computational complexity of the inference attack via interval analysis is $\mathcal{O}(n^{3})$.
-Proof 3.3.
+Proof 3.4.
 The dominant operation in the inference attack involves matrix multiplication.
@@
 3.4 Properties of the Interval Inference Attack
-The inference attack exhibits certain properties.
-For instance, the radius of the uncertainty $\mathcal{Y}_{k|k}$, i.e., $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$ is bounded by a function of the radius of the disturbance and the observation set, as stated below.
+The inference attack exhibits several key properties.
+In particular, the radius of the uncertainty set $\mathcal{Y}_{k|k}$, $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$, is bounded by a function of the radius of the disturbance and the observation set, as formalized below.
 Lemma 8.
-For any $k\geqslant 1$, the radius of $\mathcal{Y}_{k|k}$ satisfies
-$p_{k|k}^{y}\leqslant$ $\left(\left|A_{3}\right|+\left|A_{4}\right|\left|A_{2}^{-1}\right|+\left|A_{4}\right|\left|A_{2}^{-1}A_{1}\right|\right)\overline{p}^{x}$
-$+\left|A_{4}\right|\left|A_{2}^{-1}B_{1}\right|p^{w,x}_{k}+\left|B_{2}\right|p^{w,y}_{k},$ (41)
-where $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$ and $p^{w,y}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$, respectively, and $|A|$ is a matrix where each element is the absolute value of the corresponding element in $A$, i.e., $|A|=\left[|a_{i,j}|\right]$.
-Proof 3.4.
+For any $k\geq 1$, the radius of $\mathcal{Y}_{k|k}$ satisfies
+$p_{k|k}^{y}$ $\leq\Big(|A_{3}|+|A_{4}||A_{2}^{-1}|+|A_{4}||A_{2}^{-1}A_{1}|\Big)\overline{p}^{x}$
+$\quad+|A_{4}||A_{2}^{-1}B_{1}|p_{k}^{w,x}+|B_{2}|p_{k}^{w,y},$ (45)
+where $\overline{p}^{x}\geq p^{m,x}_{j|j}$ for any $j\geq 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$, and $p^{w,y}_{k}$ are the radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$, and $\mathcal{W}_{k}^{y}$, respectively, and $|A|$ denotes the matrix with elementwise absolute values, i.e., $|A|=[|a_{i,j}|]$.
+Proof 3.5.
 See Appendix B.
-Since the volume of $\mathcal{Y}_{k|k}$ is the sum of $p_{k|k}^{y}$, $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is also bounded by a function of the radius of the observation set $\mathcal{M}_{k|k}^{x}$.
-Consequently, if $\mathcal{M}_{k|k}^{x}$ is small, then the privacy level would be low, and the adversary retains little uncertainty after performing the inference attack.
-Furthermore, by comparing the predicted and posterior uncertainty sets, as given by (5) and (10), the amount of uncertainty reduction can be quantified by $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, where
-$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\setminus\mathcal{Y}_{k|k},$
-is the difference between $\mathcal{Y}_{k|k-1}$ and $\mathcal{Y}_{k|k}$.
-Since $\mathcal{Y}_{k|k}$ is a subset of $\mathcal{Y}_{k|k-1}$, $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ can be computed with
-$\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\!.$ (42)
-Therefore, to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, it is equivalent to reduce the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ since the amount of prior uncertainty $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$ is fixed at time $k$.
-As shown in the next theorem, the amount of uncertainty reduction, i.e., $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, is bounded by the new information extracted from $\mathcal{M}^{x}_{k|k}$.
+Since the volume of $\mathcal{Y}_{k|k}$ is given by the product of $2p_{k|k}^{y}$ over all dimensions, $\mathrm{Vol}(\mathcal{Y}_{k|k})$ is similarly bounded by a function of the radius of the observation set $\mathcal{M}_{k|k}^{x}$.
+Consequently, a small $\mathcal{M}_{k|k}^{x}$ implies low privacy, and the adversary retains limited uncertainty after performing the inference attack.
+Furthermore, by comparing the predicted and posterior uncertainty sets, as in (13) and (18), the reduction of uncertainty can be quantified using the surrogate measure $\overline{\mathrm{Vol}}(\Delta\mathcal{Y}_{k|k})$, where
+$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\setminus\mathcal{Y}_{k|k}.$ (46)
+Since $\mathcal{Y}_{k|k}\subseteq\mathcal{Y}_{k|k-1}$, the surrogate volume can be computed as
+$\overline{\mathrm{Vol}}(\Delta\mathcal{Y}_{k|k})=\overline{\mathrm{Vol}}(\mathcal{Y}_{k|k-1})-\overline{\mathrm{Vol}}(\mathcal{Y}_{k|k}).$ (47)
+Thus, increasing the privacy level $\overline{\mathrm{Vol}}(\mathcal{Y}_{k|k})$ is equivalent to reducing the amount of uncertainty reduction $\overline{\mathrm{Vol}}(\Delta\mathcal{Y}_{k|k})$, since the prior uncertainty $\overline{\mathrm{Vol}}(\mathcal{Y}_{k|k-1})$ is fixed at time $k$.
+As shown in the next theorem, the amount of uncertainty reduction, is bounded by the new information extracted from $\mathcal{M}^{x}_{k|k}$.
 Theorem 9.
 The amount of uncertainty reduction at $k$ is
-$\!\!\!\!\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}\oplus\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (43)
-satisfying
-$\mathrm{Vol}$ $\left(\Delta\mathcal{Y}_{k|k}\right)\geqslant 2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1},$
-$\!\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant$ $\!\left\|\!A_{3}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
-where $\Delta\mathcal{X}_{k-1|k}\!\!=\!\!\mathcal{X}_{k-1|k-1}\!\setminus\!\mathcal{X}_{k-1|k}$, $\Delta\!\mathcal{Y}_{k-1|k}\!\!=\!\!\mathcal{Y}_{k-1|k-1}\!\setminus\!\mathcal{Y}_{k-1|k}$ are the adversary’s uncertainty reduction of $X_{k-1}$ and $Y_{k-1}$, and $\left\|c_{k|k}^{y}\!-\!c_{k|k-1}^{y}\right\|_{1}$ quantifies the difference in central estimation with and without considering $\mathcal{X}_{k|k}$.
-Proof 3.5.
+$\!\!\!\!\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}\oplus\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (48)
+with bounds
+$\overline{\mathrm{Vol}}$ $\left(\Delta\mathcal{Y}_{k|k}\right)\geqslant 2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1},$
+$\!\overline{\mathrm{Vol}}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant$ $\!\left\|\!A_{3}\!\right\|_{1}\!\overline{\mathrm{Vol}}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|_{1}\!\overline{\mathrm{Vol}}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
+where $\Delta\mathcal{X}_{k-1|k}\!\!=\!\!\mathcal{X}_{k-1|k-1}\!\setminus\!\mathcal{X}_{k-1|k}$, $\Delta\!\mathcal{Y}_{k-1|k}\!\!=\!\!\mathcal{Y}_{k-1|k-1}\!\setminus\!\mathcal{Y}_{k-1|k}$, and $\|c_{k|k}^{y}-c_{k|k-1}^{y}\|_{1}$ quantifies the change in the central estimate due to the observation $\mathcal{X}_{k|k}$.
+Proof 3.6.
 See Appendix C.
-According to Theorem 9, the reduction of uncertainty $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is highly correlated with the amount of information that the adversary extracts from the observation set $\mathcal{M}_{k|k}^{x}$.
-Furthermore, with (42) and Theorem 9, we can show the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k-1|k}\right)$ is bounded in the following lemma.
+Combining (47) and Theorem 9, the privacy level $\overline{\mathrm{Vol}}(\mathcal{Y}_{k|k})$ can be bounded as follows.
 Lemma 10.
-The privacy level can be bounded with the following inequalities,
-$\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
-$\leqslant\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k}\!\right)\!\leqslant\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$
-As a result, we can reduce the extracted information $\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ via designing proper observation set $\mathcal{M}^{x}_{k|k}$.
-Moreover, when the privacy level is high, the adversary’s ability to update its central estimate $c_{k|k}^{y}$ is also limited, as indicated by the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$.
+The privacy level satisfies
+$\overline{\mathrm{Vol}}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|_{1}\!\overline{\mathrm{Vol}}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|_{1}\!\overline{\mathrm{Vol}}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
+$\leqslant\overline{\mathrm{Vol}}\!\left(\!\mathcal{Y}_{k|k}\!\right)\!\leqslant\overline{\mathrm{Vol}}\left(\mathcal{Y}_{k|k-1}\right)-2\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}.$
+As a result, we can reduce the extracted information $\overline{\mathrm{Vol}}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\overline{\mathrm{Vol}}\left(\mathcal{Y}_{k|k}\right)$ via designing proper observation set $\mathcal{M}^{x}_{k|k}$.
+Moreover, if the privacy level is high, the adversary’s ability to update its central estimate $c_{k|k}^{y}$ is also limited, as indicated by the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$.
 Based on this observation, $\mathcal{M}^{x}_{k|k}$ can also be designed to hinder accurate central estimate updates, further improving the privacy level.
@@
 Therefore, $\mathcal{X}_{k|k-1}$ is the maximum observation set $\mathcal{M}^{x}_{k|k}$ that the filter can release, i.e., $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$.
-To maintain the high data utility, the observation set has to satisfy the following constraint,
-$\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}.$
+To maintain high data utility, the uncertainty set $\mathcal{X}_{k|k}$ must satisfy the following constraint:
+$\mathrm{Vol}\left(\mathcal{X}_{k|k}\right)\leq\epsilon^{x},$
+where $\epsilon^{x}>0$ specifies the desired upper bound on the uncertainty volume.
+Since $\mathcal{X}_{k|k}$ is a subset of the observation set $\mathcal{M}_{k|k}^{x}$, this constraint can be equivalently enforced on the larger set, $\mathrm{Vol}\left(\mathcal{M}_{k|k}^{x}\right)\leq\epsilon^{x}$, which simplifies the design of the observation set while ensuring that the utility requirement is satisfied.
 To reduce privacy leakage while preserving data utility, we design the privacy filter illustrated in Fig.
@@
 […]
-Note that $\mathcal{S}^{x}_{k|k}$ is randomly generated as a subset of $\mathcal{X}_{k|k-1}$.
-In practice, it can be sufficiently small; for instance, it may contain only the true public state $x_{k}$.
-Consequently, recovering $\mathcal{S}^{x}_{k|k}$ could lead to potential privacy leakage and should therefore be avoided.
-In the following, we show that an attacker cannot recover $\mathcal{S}^{x}_{k|k}$ by inverting the optimization problem $\mathbf{P_{1}}$ due to the randomization operation.
+Note that $\mathcal{S}^{x}_{k|k}$ is randomly generated as a subset of $\mathcal{X}_{k|k-1}$ and can be made sufficiently small in practice.
+For instance, the random set $\mathcal{S}^{x}_{k|k}$ may contain only the true public state $x_{k}$.
+According to Lemma 8, a sufficiently small observation set could lead to potential privacy leakage.
+Therefore, recovering $\mathcal{S}^{x}_{k|k}$ from $\mathbf{P_{1}}$ must be avoided.
+In the following, we demonstrate that an attacker cannot recover $\mathcal{S}^{x}_{k|k}$ by inverting the optimization problem $\mathbf{P_{1}}$, owing to the randomization mechanism embedded in the privacy filter.
 Proposition 11.
@@
 (2) In the presence of the inference attack described in (14)–(20), the filter output maximizes the privacy level.
-(3) The filter is robust against reverse attacks that attempt to recover the sufficiently small set $\mathcal{S}_{k|k}^{x}$.
-Moreover, since the computational complexity of both the inference attack in (6)–(12) and the volume computation increases with the complexity of the set representations, a trade-off exists between the achievable privacy enhancement of the proposed filter and its computational cost.
-While more sophisticated set representations could improve the accuracy of privacy evaluation, for efficiency and clarity, we next present a concrete design based on the interval approximation described in Sec.
-3.2.
+(3) The filter is robust against reverse attacks that attempt to recover the sufficiently small set $\mathcal{S}_{k|k}^{x}$, thereby reducing the risk of privacy compromise from adversaries exploiting structural vulnerabilities.
+Moreover, the computational complexity of both the inference attack in (14)–(20) and the volume computation increases with the complexity of the set representations.
+Consequently, a trade-off exists between the achievable privacy enhancement of the proposed filter and its computational cost.
+While more sophisticated set representations may improve the accuracy of privacy evaluation, for efficiency and clarity, we next present a concrete design based on the interval approximation described in Sec.
+3.3.
+In addition, as discussed in Sec.3.3, the surrogate measure $\overline{\mathrm{Vol}}(\cdot)$ bounds the volume of interval, and could simplify the computation complexity.
+We adopt the surrogate volumetric measure $\overline{\mathrm{Vol}}(\cdot)$ in the optimization to further reduce computational complexity, and we verify that using this surrogate measure still leads to improved privacy levels in Sec.
+5.
 Figure 2: Structure of the proposed privacy filter.
@@
 Also, we can shown $\mathcal{S}^{x}_{k|k}$ satisfies the utility constraint as follows,
-$\mathrm{Vol}\left(\mathcal{S}^{x}_{k|k}\right)=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
+$\overline{\mathrm{Vol}}\left(\mathcal{S}^{x}_{k|k}\right)=$ ${\beta}_{k}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}+{\alpha}_{k}\left\|x_{k}-\underline{X}_{k|k-1}\right\|_{1}$
 $\leqslant$ $\frac{\epsilon^{x}}{2\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}}\left\|\overline{X}_{k|k-1}-x_{k}\right\|_{1}$
@@
 Theorem 12.
-The privacy filter optimization problem $\mathbf{P_{1}}$ is equivalent to the following linear programming
+The privacy filter optimization problem $\mathbf{P_{1}}$ with the surrogate privacy measure $\overline{\mathrm{Vol}}\left(\mathcal{Y}_{k|k}\right)$ and utility measure $\overline{\mathrm{Vol}}\left(\mathcal{M}^{x}_{k|k}\right)$ is equivalent to the following linear programming
 $\mathbf{P_{2}}:$ $\max_{\epsilon^{y},\mathcal{M}^{x}_{k|k},p_{k-1|k}^{\Delta x},p_{k-1|k}^{\Delta y}}\epsilon^{y}$
@@
 Figure 3: Inference attack results after applying the optimal privacy filter: (a), (b) Estimated $X_{k}$ and (e), (f) Estimated $Y_{k}$ for $\mathrm{Vol}\!\left(\mathcal{M}_{k|k}^{x}\right)\leq 0.01$; (c), (d) Estimated $X_{k}$ and (g), (h) Estimated $Y_{k}$ for $\mathrm{Vol}\!\left(\mathcal{M}_{k|k}^{x}\right)\leq 0.5$.
-Consequently, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 2.3.
+Consequently, we can solve the linear programming problem $\mathbf{P_{2}}$ to obtain the optimal observation set that defends the system against the inference attack defined in Section 3.
 5 Numerical Verification
@@
 As discussed in Section 2, the production rate is private but the inventory information has to be released.
-We first plot the trajectories of system states and their interval tubes in Fig.3 under the optimal privacy filter design for different values of $\epsilon^{x}$.
-We also use the central point of the posterior intervals as one of the possible testing estimation, e.g., $\frac{\overline{X}_{k|k}+\underline{X}_{k|k}}{2}$ for $\mathcal{X}_{k|k}$.
-The pink areas in these figures represent the uncertainty sets of system states.
-As shown in Fig.3, when $\epsilon^{x}=0.01$ is small, the adversary’s uncertainty about the private production rate is small, and its central estimation closely matches the actual production rate.
-However, when $\epsilon^{x}$ increases to $0.5$, the utility of the inventory information decreases slightly, but this leads to higher uncertainty of the inference attack, causing the adversary’s central estimation of the production rate to become less accurate.
-This observation numerically verifies Theorem 9, confirming that a higher privacy level prevents the adversary from refining its incorrect central estimation.
-Therefore, the proposed privacy filter effectively reduces the privacy leakage of the production rate, though at the cost of introducing some inaccuracies in the inventory information.
+We first plot the trajectories of the system states and their corresponding interval tubes in Fig.
+3 under the optimal privacy filter design for different values of $\epsilon^{x}$.
+The shaded pink areas represent the interval tubes, i.e., the uncertainty sets of the system states, which quantify the range of values that an adversary can infer.
+As shown in Fig.
+3, for $\epsilon^{x}=0.01$, the adversary’s uncertainty about the private production rate is small.
+However, increasing $\epsilon^{x}$ to $0.5$ slightly reduces the utility of the inventory information but significantly enlarges the adversary’s uncertainty.
+To further illustrate the effectiveness of the privacy filter, we consider one possible adversary estimate based on the central points of the posterior intervals.
+For the public state $\mathcal{X}_{k|k}$, the adversary’s central estimate is given by $\frac{\overline{X}_{k|k}+\underline{X}_{k|k}}{2}$.
+As shown in Fig.
+3, when $\epsilon^{x}$ increases from $0.01$ to $0.5$, the adversary’s central estimate of the production rate becomes significantly less accurate, whereas the central estimate of the inventory information remains accurate.
+This observation numerically confirms Theorem 9, showing that a higher privacy level prevents the adversary from refining an incorrect central estimate.
+Hence, the proposed privacy filter effectively mitigates the leakage of production rate information, while introducing a controlled loss of accuracy in the inventory information.
 Figure 4: The truncated Gaussian mechanism.
@@
 Because the additive noise $v_{k}$ is bounded within $\left[-\epsilon^{x}/2,\,\epsilon^{x}/2\right]$, the publicly released state is guaranteed to lie within the interval $\mathcal{M}_{k|k}^{x}$.
-The privacy-utility trade-off was evaluated by plotting the average privacy level of the production rate against the average utility of the inventory over 100 random trajectories with horizon 100, as illustrated in Fig.
+We evaluate the privacy-utility trade-off by plotting the average privacy level of the production rate against the average utility of the inventory in Fig.
 5.
-For a fair and clear comparison, the privacy level and utility values for the truncated Gaussian mechanism are normalized to the range $[0,1]$, and its same scaling parameters are applied to the other two mechanisms.
+For a fair and clear comparison, the privacy level and utility values for the truncated Gaussian mechanism are normalized to the range $[0,1]$, and the same scaling parameters are applied to the other two mechanisms.
 The results demonstrate that an increase in inventory utility corresponds to a reduction in the privacy level of the production rate, thereby confirming the intrinsic trade-off between data utility and privacy protection.
-Besides, the quantization mechanism and the truncated Gaussian mechanism have similar performance in privacy-utility trade-off.
-Note that these two mechanisms assume static domains of states, while the sets of states in systems subject to disturbance can be estimated and described more precise as addressed by the inference attack.
-The proposed volumetric approach estimates the time-varying private state set and adjusts the output accordingly.
-Furthermore, as discussed in [31], if the volume (e.g., interval length) of domain satisfies certain conditions, the truncated Gaussian mechanism ensures differential privacy.
+As discussed in [31], if the interval length, i.e., volume, of domain satisfies certain conditions, the truncated Gaussian mechanism ensures differential privacy.
 However, given the volume constraint, the shape of the public state set can be arbitrary, and certain shapes may lead to substantial volumetric leakage of the private state after inference attacks.
@@
 In this paper, we develop a volumetric framework for privacy analysis and defense in dynamic systems subject to bounded disturbance.
-An inference attack, whereby an adversary estimates the private state, is formalized, and a volumetric measure is introduced to quantify the resulting privacy level.
+An inference attack, whereby an adversary estimates the private information, is formalized, and a volumetric measure is introduced to quantify the resulting privacy level.
 We develop computational methods based on interval analysis, and establish the theoretical properties of the measure.
@@
 which satisfies $p_{k-1|k}^{\Delta z}\geqslant 0$. As a result, we have
-$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ $=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
+$\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k|k}\right)$ $=\left\|\left|A_{3}\right|p_{k-1|k}^{\Delta x}+\left|A_{4}\right|p_{k-1|k}^{\Delta y}\right\|_{1}$
 $\overset{(a)}{\leqslant}\left\|A_{3}\right\|_{1}\left\|p_{k-1|k}^{\Delta x}\right\|_{1}+\left\|A_{4}\right\|_{1}\left\|p_{k-1|k}^{\Delta y}\right\|_{1}$
-$=\left\|A_{3}\right\|_{1}\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|_{1}\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right),$
+$=\left\|A_{3}\right\|_{1}\overline{\mathrm{Vol}}\left(\Delta\mathcal{X}_{k-1|k}\right)+\left\|A_{4}\right\|_{1}\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k-1|k}\right),$
 where $(a)$ is due to $p_{k-1|k}^{\Delta x}\geqslant 0$ and $p_{k-1|k}^{\Delta y}\geqslant 0$.
@@
 Therefore, we have
-$\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
+$\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k|k}\right)=$ $\left\|\overline{Y}_{k|k}-\overline{Y}_{k|k-1}\right\|_{1}+\left\|\underline{Y}_{k|k}-\underline{Y}_{k|k-1}\right\|_{1}$
 $\geqslant$ $\left\|\overline{Y}_{k|k}+\underline{Y}_{k|k}-\left(\overline{Y}_{k|k-1}+\underline{Y}_{k|k-1}\right)\right\|_{1}$
@@
 Therefore, we can replace the objective function with the slack variable $\epsilon^{y}$ and add $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\leqslant\epsilon^{y}$ as a new constraint, and then minimize $\epsilon^{y}$.
-Since $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is determined by $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, we can replace constraints (28) and (31) with the constraints of difference sets (C).
+Since $\overline{\mathrm{Vol}}\left(\Delta\mathcal{Y}_{k|k}\right)$ is determined by $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, we can replace constraints (32) and (35) with the constraints of difference sets (C).
 Also, the objective function increases with any elements of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ since the elements of $p_{k-1|k}^{\Delta x}$, $p_{k-1|k}^{\Delta y}$, $\left|A_{3}\right|$ and $\left|A_{4}\right|$ are non-negative.
 As a result, we can replace the constraint of $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$, i.e., (C), with inequalities (68), and let $p_{k-1|k}^{\Delta x}$ and $p_{k-1|k}^{\Delta y}$ be decision variables.
-Besides, the constraints $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ and $\overline{M}^{x}_{k|k}\geqslant\underline{M}^{x}_{k|k}$ are equivalent to the inequality constraint, $\underline{X}_{k|k-1}\leqslant\underline{M}^{x}_{k|k}\leqslant\underline{S}^{x}_{k|k}\leqslant\overline{S}^{x}_{k|k}\leqslant\overline{M}^{x}_{k|k}\leqslant\overline{X}_{k|k-1}$, and the utility constraint $\mathrm{Vol}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ can be replaced with $\left\|\overline{M}^{x}_{k|k}-\underline{M}^{x}_{k|k}\right\|_{1}\leqslant\epsilon^{x}$.
+Besides, the constraints $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$, $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ and $\overline{M}^{x}_{k|k}\geqslant\underline{M}^{x}_{k|k}$ are equivalent to the inequality constraint, $\underline{X}_{k|k-1}\leqslant\underline{M}^{x}_{k|k}\leqslant\underline{S}^{x}_{k|k}\leqslant\overline{S}^{x}_{k|k}\leqslant\overline{M}^{x}_{k|k}\leqslant\overline{X}_{k|k-1}$, and the utility constraint $\overline{\mathrm{Vol}}\left(\mathcal{M}^{x}_{k|k}\right)\leqslant\epsilon^{x}$ can be replaced with $\left\|\overline{M}^{x}_{k|k}-\underline{M}^{x}_{k|k}\right\|_{1}\leqslant\epsilon^{x}$.
 Finally, the objective and the constraints are linear functions of the decision variables, thus, the optimal privacy filter can be obtained by solving the linear programming $\mathbf{P_{2}}$.